### v2

# Reconstructing Hand-Held Objects in 3D

Abstract
Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from in-the-wild RGB images or videos. Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels. At the same time, two strong anchors emerge in this setting: (1) estimated 3D hands help disambiguate the location and scale of the object, and (2) the set of manipulanda is small relative to all possible objects. With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets. Our model, MCC-Hand-Object (MCC-HO), jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs. Subsequently, we use GPT-4(V) to retrieve a 3D object model that matches the object in the image and rigidly align the model to the network-inferred geometry; we call this alignment Retrieval-Augmented Reconstruction (RAR). Experiments demonstrate that MCC-HO achieves state-of-the-art performance on lab and Internet datasets, and we show how RAR can be used to automatically obtain 3D labels for in-the-wild images of hand-object interactions.
1 Introduction
“Man is the measure of all things.”
Protagoras
In manipulation, one may philosophize that the hand is the measure of all objects. From infancy, humans use their hands to gather information about the physical world [30, 83, 64, 67]. Through our hands and eyes, we learn to estimate 3D geometry and physical properties as well as affordances that inform the ways in which we manipulate objects [39, 14].
Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics. While there has been significant progress in reconstructing hands [63, 50, 52] and objects [85, 37, 36] separately, joint reconstruction of hand-object interactions presents unique challenges that cannot currently be solved by general reconstruction approaches. Procedural methods such as Structure from Motion (SfM) [72] will struggle because the object is mostly occluded, and thus it is difficult to obtain enough feature points to achieve accurate reconstruction. Similarly, Neural Radiance Fields (NeRFs) [44] are best suited for unoccluded multiview reconstruction of static scenes and objects. Instead, state-of-the-art approaches in hand-object reconstruction have relied on human assistance for data collection [3, 51, 13]. While manually-retrieved 3D object models naturally improve reconstruction quality and can provide better initialization, this paradigm is not scalable. This is because human-driven retrieval typically includes searching online for a suitable object model, verifying that the model matches the object in the image, and providing an estimate of 3D object pose [3]. Rigidly aligning the 3D model to an image is also an underconstrained 3D-to-2D problem, and optimization approaches are prone to issues with local minima.
Figure 1: We present a scalable approach to hand-held object reconstruction that is guided by recognition and retrieval. In the first stage, our transformer-based model (MCC-HO) infers hand-object 3D geometry from a single RGB image (Row 2, visualized as meshes). In the second stage, a template object model is obtained using Retrieval-Augmented Reconstruction (RAR). This template model is then rigidly aligned with the output of MCC-HO using ICP (Row 3).
In this paper, we present a novel approach to reconstructing hands manipulating common objects from monocular RGB image(s) or video. We draw from three key insights:
  1. 1.
    One can recover shape even from single images at test time, if the object being seen has some family resemblance to previously seen 3D objects. For children, the ability to see 3D from motion comes first, followed by binocular stereopsis, and then finally they can guess 3D shape even from a single image [92]. An operationalization of this in computer vision is the MCC system [85] — the model requires 3D information at training time, but at test time a single image is enough.
  2. 2.
    In images of hand-held objects, we expect to see hands. The hands also provide a strong anchor for 3D location and scale; for instance, the hands provide a convenient measurement tool for many commonplace activities (e.g., cooking, crafts). Recent progress in single-image hand reconstruction [52] enables the development of a robust hand-centric approach to object reconstruction.
  3. 3.
    Multiple datasets have now been collected which have videos of hand-object interactions with underlying 3D ground truth (see Section 2), but there has not been a model trained on all available data.
Given an RGB image and estimated 3D hand, we train a transformer-based model, MCC-Hand-Object (MCC-HO), that jointly infers 3D hand and object geometry. This geometry is represented as a neural implicit surface composed of occupancy, color, and hand-object segmentation. Our model architecture is adapted from an existing object reconstruction model, MCC [85], and thus we make use of a learned object prior by fine-tuning an MCC model that is trained on the CO3Dv2 dataset [59].
In conjunction with MCC-HO, we propose Retrieval-Augmented Reconstruction (RAR), a method for automatically retrieving object models using large language/vision models to improve network-inferred geometry. Specifically, we prompt GPT-4(V) [48] to recognize the hand-held object in the image and provide a detailed text description. This description is passed to Genie [42], a text-to-3D generative model, to obtain a 3D object. After the object is retrieved, iterative closest point (ICP) [1] or other point cloud fitting techniques can be used to rigidly align the Genie mesh with the estimated geometry from our network. See Figure 1.
To summarize, this paper presents a paradigm for reconstructing hand-held objects that combines the respective advantages of model-free and model-based approaches. Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment of the retrieved model with network-inferred geometry. In the first stage, MCC-HO jointly estimates hand and object geometry from a single RGB image and 3D hand. In the second stage, RAR is used to improve reconstruction quality via recognition and retrieval. The hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [42]). In the third stage, the retrieved 3D object is rigidly aligned with the network-inferred object geometry. Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [5, 3, 38]. To demonstrate the scalability of our approach, we additionally use RAR to obtain ground truth 3D hands and objects for previously unlabeled images in the 100DOH dataset [66].
2 Related Work
Object reconstruction. Reconstructing objects from images is a longstanding problem in computer vision, and the bulk of research in this area has focused on scenarios where the object is fully or almost fully unoccluded in the input view(s). Structure from motion (SfM) [21, 65, 68, 72] and SLAM [4, 47, 69] remain robust solutions particularly for multiview reconstruction and camera parameter estimation. Learning-based object reconstruction methods typically represent 3D geometry as point clouds [54, 55], meshes [16, 81, 84], voxel grids [86, 78], or implicit surfaces [49, 44, 46, 71, 79, 85]. Following the release of Objaverse [11] and Objaverse-XL [10], a number of works have used this large-scale 3D object dataset to train monocular object reconstruction models [37, 36, 40, 26]. The CO3Dv2 dataset [60] contains objects from 50 MS-COCO categories [35] and has also been used to train monocular object reconstruction models [85]. In practice, we find that off-the-shelf object reconstruction methods do not perform well on images of hand-object interactions due to occlusion from the hand, larger object pose diversity, etc.
Hand reconstruction. Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [52, 34, 43, 45, 76, 93, 96, 50, 63] or monocular video [28, 82, 88]. These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [62]. Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [77, 94]. Particularly relevant to our work, [61] considers the challenging case of predicting pose when the hand is interacting with objects, which can lead to large occlusions of the hand and/or object.
Model-based hand-object pose estimation. When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [2, 3, 51, 13, 22, 23, 70, 74, 89, 87, 73, 56]. The estimated hands can then be used to rigidly align the object with images or videos [3, 51] via inverse rendering [41, 58]. However, image-based optimization losses are prone to issues with local minima, as depth/3D information is ambiguous from a single RGB image.
Model-free hand-object reconstruction. Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions. Existing approaches are either tailored to reconstruction from a single image [90, 24, 7, 6, 95, 15, 8], RGB video [27, 91, 19, 53], or RGB-D video [80]. Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved template 3D model.
Hand-object datasets. Datasets containing hand-object interactions are either collected in lab environments [2, 20, 5, 31, 38], through large-scale capture [17, 18], or from Internet videos [9, 66, 3]. A number of hand-held object reconstruction models [7, 6, 8] were trained on either the HO-3D [20] or DexYCB [5] datasets. Since the combined number of object instances in these two datasets is 31, it is difficult for such approaches to generalize to unseen objects. HOI4D [38] is currently the largest available dataset with 3D hand and object labels (also used in [91]). There is also a comparably larger amount of data containing videos of hand-object interactions without 3D labels. Epic Kitchens [9] contains 55 hours of egocentric video of people in their personal kitchens. Ego4D [17] and Ego-Exo4D [18] are also egocentric datasets, and 100 Days of Hands (100DOH) [66] contains 131 days of footage collected from Internet videos. Following the release of [66], MOW [3] and RHOV [51] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
Figure 2: Given an RGB image and estimated 3D hand, our method reconstructs hand-held objects in three stages. In the first stage, MCC-HO is used to predict hand and object point clouds. In the second stage, a template mesh for the object is obtained using Retrieval-Augmented Reconstruction (RAR). In the third stage, the template object mesh is rigidly aligned to the network-inferred geometry using ICP.
3 Method Overview
We propose a three-stage approach to reconstructing hand-object interactions that combines feedforward network inference with object model retrieval and subsequent 3D fitting. See Figure 2. Our approach is built on two important insights in the hand-object interaction setting. The presence of a human hand provides grounding in terms of both the 3D location and scale of the object. While monocular hand reconstruction is an ill-posed problem and can be a limiting factor for hand-conditioned object reconstruction approaches, recent progress has enabled accurate hand estimation by scaling up [52]. Thus, as in prior work [90, 8], the estimated object pose and shape is conditioned on the estimated 3D hand. Applying a hand-centric approach has several advantages. Since monocular image/video reconstruction is particularly ill-posed with respect to depth and scale, 3D hands can be used to disambiguate the location and size of the object. Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
In addition, although the hand introduces significant occlusions of the object, the space of possible manipulanda is much smaller than the space of all objects. Thus, this work focuses on reconstructing categories of common household objects, similar in spirit to how the COCO dataset [35] was constructed and in alignment with object [60] and hand-object [38] datasets. The pervasiveness of such commonplace objects enables deploying large image recognition systems for the task of automatically recognizing the object and retrieving a 3D object model. Since household manipulanda are often quite uniform in shape (e.g., a knife, an 8 o.z. mug), the accuracy of such a retrieval process is much higher than for generic objects, whether through lookup in an object dataset [11, 10] or text-to-3D generation methods [42] trained on such data.
4 Hand-Object Geometry Inference
In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [85] for the specific task of hand-object reconstruction. Notably, two domain shifts from the original MCC model are applied. First, the input depth image is removed and the input RGB-D point cloud is replaced with the 3D hand. Second, the model is trained to infer both hand and object geometry, and thus an extra segmentation output is added to the decoder. We call our model MCC-Hand-Object (MCC-HO).
The inputs to MCC-HO are a single RGB image and 3D hand geometry. Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM) [29], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image. Hand geometry is typically parameterized by the MANO hand model [62], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$. The MANO hand skeleton contains 21 joints. At training time, the ground truth 3D hand is used. At test time, HaMeR [52] or any other 3D hand estimation methods can be used.
The inferred geometry is represented as a neural implicit function $\rho(x)$ modeling the hand and object jointly, such that for any 3D location $x$, $\rho$ returns the occupancy probability $\sigma(x)$, RGB color $c(x)$, and segmentation label $m(x)$:
$\rho(x)=(\sigma(x),c(x),m(x))$ (1)
where $\sigma(x)\in[0,1]$, $c(x)\in[0,1]^{3}$, and $m(x)\in\{0,1,2\}$ to indicating a background, hand, or object point. The scale of the object geometry is normalized with respect to the input 3D hand.
4.1 Architecture
MCC-HO has an encoder-decoder architecture that is composed of transformers [75] and outputs an implicit representation of hand-object geometry. The input images and 3D hand are passed to the encoder to compute a feature map that is a function of both image and hand embeddings. These features are then used to condition the decoder, which maps any 3D location to occupancy, color, and hand-object segmentation. During training, a set of query 3D points $Q$ are uniformly sampled within the bounds of a normalized volume. At test time, vertices of a voxel grid are passed as query points.
Encoder. The encoder takes as input a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a posed MANO hand mesh defined in camera world space with vertices $v_{h}\in\mathbb{R}^{778\times 3}$ and faces $f_{h}\in\mathbb{R}^{1538\times 3}$. The camera intrinsics are either known or consistent with the estimated 3D hand (e.g., the camera parameters used during HaMeR inference [52]). In order to jointly encode the image and hand geometry, the hand mesh is rasterized and per-pixel 3D points for the hand are sampled. For each pixel $(i,j)$ in the image, a ray is cast from the camera aperture to the pixel center to determine the first intersection point with the hand mesh at a point $P_{0}(i,j)$ in camera world space. If an intersection exists, the 3D location of each point is barycentrically interpolated from the intersecting camera world space triangle vertices, denoted $\{v_{h,1},v_{h,2},v_{h,3}\}$, such that:
$P_{0}(i,j)=\alpha_{1}v_{h,1}+\alpha_{2}v_{h,2}+\alpha_{3}v_{h,3}\in\mathbb{R}^{3},$ (2)
with barycentric weights $\alpha$. The points $P_{0}$ are used to define a normalized camera world space, such that the 3D coordinates passed to the network are $P=s(P_{0}-\mu(P_{0}))/\sigma(P_{0})$ and geometry is inferred in this same coordinate system. Importantly, $s$ is a constant scaling factor used to ensure the object geometry can be contained by the bounds of this hand-normalized sampling space. See Figure 3.
Figure 3: The input 3D hand is rasterized and per-pixel 3D points are sampled to be passed to the MCC-HO encoder.
The image $I$ and hand points $P$ are encoded into a single representation $R$, where:
$R\coloneqq f(E^{RGB}(I),E^{XYZ}(P))\in\mathbb{R}^{N^{enc}\times C},$ (3)
and $E^{RGB}$ and $E^{XYZ}$ are identical to the transformers proposed in MCC [85], though we define $P$ differently, i.e., without an input depth image. The image transformer $E^{RGB}$ has a Vision Transformer (ViT) architecture [12] with $16\times 16$ patch embeddings. The point transformer $E^{XYZ}$ uses a self-attention-based patch embedding design that differentiates between seen and unseen pixel points. In particular, for points in $P$ where no ray-mesh intersection was found, a special $C$-dimensional embedding is learned in lieu of embedding 3D points. The function $f$ concatenates the output of the two transformers, each of which has dimensions $N^{enc}\times C$. $N^{enc}$ is the number of tokens used in the transformers and $C$ is the output channel size.
Decoder. The decoder takes as input the encoder output, $R$, as well as a set of query points $Q\in\mathbb{R}^{N_{q}\times 3}$ (we choose $N_{q}=1024$). These inputs are passed to a transformer with architecture adopted from MCC and inspired by Masked Autoencoders (MAE) [25]. Each query token is passed to three output heads that infer binary occupancy $\sigma$, RGB color $c$, and hand-object segmentation $m$, such that:
$Decoder(R,Q)\coloneqq(\sigma(Q),c(Q),m(Q)).$ (4)
The decoder layers that output occupancy and color align with MCC, but MCC-HO has an additional linear layer that infers multiclass segmentation of each point into either background, hand, or object labels.
As discussed in further detail in MCC, a point-query-based decoder presents a number of advantages with respect to computational efficiency and scalability. Since the encoder is independent of the point queries, any number of points can be sampled using the same encoder structure and feature map size. The decoder architecture is relatively lightweight and also enables dynamic changes to the point sampling resolution and query size at training and/or test time.
4.2 Implementation Details
The training loss is a combination of occupancy, color, and segmentation:
$\mathcal{L}=\mathcal{L}_{\sigma}+\mathcal{L}_{c}(\sigma_{gt})+\mathcal{L}_{m}$ (5)
The occupancy loss $\mathcal{L}_{\sigma}$ is a binary cross-entropy loss comparing the predicted and ground truth occupancies of each query point. The color loss $\mathcal{L}_{c}$ is a 256-way classification loss computed for query points that have ground truth occupancy of 1. The segmentation loss $\mathcal{L}_{m}$ is 3-way classification loss for each query point corresponding to background (empty space), hand, and object labels.
Our model was trained on a single Quandro RTX 8000 for 3k epochs with a learning rate of $1e^{-5}$ and batch size of 16. We use a cosine schedule and a linear warm-up for the first 5 epochs. At each epoch, one random frame from each sequence is selected. Network parameters that correspond to the MCC architecture [85] were initialized using a pretrained MCC model that was trained on the CO3Dv2 dataset [60]. The weights of the additional linear layer in our decoder that outputs hand-object segmentation are initialized via Xavier initialization.
5 Retrieval-Augmented Reconstruction
In order to further improve upon the network-inferred object geometry, we use GPT-4(V) [48] to automatically detect and retrieve a 3D object corresponding to the hand-held object in an image. We coin this technique Retrieval-Augmented Reconstruction (RAR) in reference to Retrieval-Augmented Generation in natural language processing [32].
Given an input image (one could also use stacked video frames), we prompt GPT-4(V) to provide a detailed description of the object the hand is holding. This text caption is passed to Genie, a text-to-3D model from Luma [42], to retrieve realistic 3D object geometry and appearance (Genie is also used in [33]). See Figure 4. After the object model is obtained, it is rigidly aligned with our network-inferred object geometry using ICP [1]. In particular, we sample a set of random initial rotations for the retrieved 3D model and select the ICP transformation that results in the lowest alignment error. Genie outputs geometry normalized with respect to a unit cube, so the scale of the 3D model is initialized using the bounding volume of the predicted object point cloud.
Figure 4: Retrieval-Augmented Reconstruction (RAR). Given an input image, we prompt GPT-4(V) [48] to recognize and provide a text description of the hand-held object. The text description is passed to a text-to-3D model (in our case, Genie [42]) to obtain a 3D object.
6 Experiments
Section 6.1 details the datasets that were used for training and evaluation. Section 6.2 defines the evaluation metrics used for comparisons. Section 6.3 quantifies the advantages of using MCC-HO, e.g., a hand-centric approach, in comparison to the original MCC model [85]. We also quantitatively (Section 6.4) and qualitatively (Section 6.5) evaluate our model against existing work in model-free, monocular hand-object reconstruction [24, 90] and find that MCC-HO achieves state-of-the-art performance on all three datasets used [5, 3, 38]. In particular, the methods most similar to ours are HO [24] and IHOI [90]. The quality of the MCC-HO predicted geometry is also shown to be sufficient for accurate rigid alignment of template objects using ICP. Also in Section 6.5.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach on in-the-wild Internet images to demonstrate the feasibility and scalability of our automated model retrieval and fitting. Specifically, we apply MCC-HO and RAR to unlabeled images from the 100DOH dataset [66].
6.1 Datasets
We experiment on four different datasets for hand-object reconstruction.
HOI4D: The HOI4D dataset [38] consists of 4,000 lab videos of hands interacting with specified categories of common objects. We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets. In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [91]. For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
DexYCB: The DexYCB dataset [5] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences (consistent with prior work [7, 6]).
MOW/RHOV: MOW [3] is a dataset of 512 images from 100 Days of Hands [66] labeled with 3D hands and objects. We use the same splits as [90] with 350 training images and 92 test images. Additionally, we use 110 videos from RHOV [51], each of which is a 2 second clip centered at the corresponding MOW frame. In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
100DOH: The 100 Days of Hands dataset [66] contains 131 days of footage collected from Internet videos showing hands interacting with objects. The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
For DexYCB [5], MOW [3], and HOI4D [38], all results generated using our approach use the ground truth 3D hand provided by each dataset. For 100DOH [66] results (e.g., examples not contained in MOW), we use HaMeR [52] to estimate the 3D hand.
6.2 Evaluation Metrics
Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics, Chamfer Distance (CD) and F-score reported at 5mm and 10mm. Both metrics are computed on point sets, and thus the point clouds inferred by MCC-HO are directly used. The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [63] and HOI4D optimizes hand pose using 2D keypoint annotations. For each example, we sample 10k points on the surface of the ground truth mesh consistent with [90, 91] (no transformations are applied to the ground truth geometry). The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well. In order to account for unknown ground truth camera parameters, ICP with scaling [58] is used to align all predicted geometry with the ground truth mesh. Median metrics are reported for each test dataset.
Chamfer Distance: The Chamfer Distance metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
$CD(A,B)=\frac{1}{N}\sum_{a\in A}\min_{b\in B}||a-b||^{2}_{2}+\frac{1}{N}\sum_{b\in B}\min_{a\in A}||a-b||^{2}_{2},$ (6)
[…]
F-score: For 3D evaluation of predicted points, the F-score at a given threshold $t$ is defined as:
$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{\text{Precision}(X_{pd},X_{gt},t)\cdot\text{Recall}(X_{pd},X_{gt},t)}{\text{Precision}(X_{pd},X_{gt},t)+\text{Recall}(X_{pd},X_{gt},t)}.$ (7)
6.3 MCC Baseline
Because our model is built on top of MCC [85], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction. For MCC, we use the released model that was trained on the CO3Dv2 dataset [60] and apply an existing monocular depth estimator [57] to obtain depth maps (as MCC takes an RGB-D image as input). Table 1 compares 3D metrics between MCC and MCC-HO computed on the MOW test set. We find that fine-tuning MCC for the specific task of hand-held object reconstruction improves performance on test images unseen by both models.
Model Parameters F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
MCC [85] 0.10 0.21 31.6
MCC-HO 0.15 0.31 15.2
[…]
DexYCB MOW HOI4D
F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
HO [24] 0.24 0.48 4.76 0.03 0.06 49.8 0.28 0.51 3.86
IHOI [90] - - - 0.13 0.24 23.1 0.42 0.70 2.7
MCC-HO 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
[…]
6.4 Quantitative Evaluation
MCC-HO is quantitatively evaluated on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [24, 90]. See Table 2. In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid. For DexYCB, we use every 5th frame of the “s0” test split, and for HOI4D, we use all frames from the 12 test sequences that contain a ground truth 3D hand label. The results for IHOI [90] and HO [24] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [91] papers. Because AlignSDF [7] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics.
Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset. We conjecture that increasing the number of unique object instances seen during training is important to model generalization. In addition, the pretrained MCC model provides a strong object prior for initialization.
6.5 Qualitative Evaluation
We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well template objects can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks). Figure 5 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud using ICP. Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera. Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact. Additionally, the object in contact with the hand is correctly identified by MCC-HO despite being provided a hand-object bounding box and no segmentation maps.
Figure 5: MCC-HO results for MOW test examples. The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud using ICP (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
Figure 6: Our approach applied to test images from MOW (Rows 1-3), DexYCB (Row 4), and HOI4D (Row 5). Each example includes the input image (left), rigid alignment of the ground truth mesh with the MCC-HO predicted point cloud using ICP (middle), and an alternative view (right).
Figure 7: Our approach applied to images in the 100 Days of Hands dataset [66]. For each image, we estimate 3D hands using HaMeR [52], inference object geometry using MCC-HO, retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie), and rigidly align the 3D model with the inferred point cloud from MCC-HO.
Figure 6 shows additional visualizations of applying rigid alignment of the template model to MCC-HO results. We include held-out test examples from all three datasets used for training: MOW, DexYCB, and HOI4D. Because MCC-HO is fine-tuned from a general object reconstruction transformer model (MCC), it is robust to unseen object instances at test time. For instance, objects in the MOW test images were never seen by MCC-HO during training. From a practical standpoint, we find 3D-to-3D fitting of template objects with the MCC-HO predicted geometry (e.g., via ICP) to be significantly more efficient than prior 3D-to-2D or per-example optimization approaches [3, 51, 91]. Moreover, optimizing 3D pose with respect to 2D images tends to get stuck in local minima. As Figure 6 demonstrates, this 3D fitting results in physically plausible hand-object interactions that when projected align with the image data.
6.5.1 Retrieval-Augmented Reconstruction.
The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images. Specifically, we take 9 images from 100DOH [66] that do not have 3D ground truth (e.g., not in MOW) and apply MCC-HO combined with RAR to reconstruct hand-object 3D geometry. See Figure 7. Estimated 3D hands for each image are obtained using HaMeR [52]. At test time, our method does not require an input hand-object segmentation mask [3, 51], depth image [85], or hand joint locations [90, 8]. Rather, we simply run feed-forward inference using MCC-HO, retrieve a 3D object mesh using RAR, and run ICP to align this mesh with the inferred point cloud. As more training data is collected through RAR and human-in-the-loop verification, we expect MCC-HO to naturally improve as well (similar in spirit to the data collection process used in Segment Anything [29]).
6.6 Limitations
While our 3D-to-3D rigid alignment approach using ICP is efficient, some misalignment may occur when rendering the reconstruction from the input view camera. This is because we do not use image-based cues for alignment, but note that this can be easily fixed via inverse rendering as in [3, 51]. In particular, the 6DOF object pose can be fine-tuned by minimizing the silhouette loss and hand-object contact/collision losses defined in [3]. See Figure 8. Regardless, our reconstruction provides a more automatic/robust initialization of the 3D object model than manual or random initialization, as was done in prior work [3, 51]. In addition, because MCC-HO is conditioned on an input 3D hand, the accuracy of the predicted hand pose naturally affects object reconstruction performance.
Figure 8: The input image is not used when running ICP, which can sometimes result in image misalignment errors (Row 1). A postprocess can be applied to resolve silhouette misalignment and any hand-object contact issues (Row 2).
7 Conclusion
In this paper, we present a novel paradigm for reconstructing hand-held objects that combines the respective benefits of model-free and model-based prediction. Our transformer-based model, MCC-HO, is trained to predict hand-object geometry given a single RGB image and estimated 3D hand. Experiments conducted using the DexYCB, MOW, and HOI4D datasets all demonstrate that MCC-HO achieves state-of-the-art performance in hand-held object reconstruction. We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets. These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH images [66]. The combination of MCC-HO and RAR can enable the construction of large hand-object datasets at-scale; for future work, we plan to expand on this paradigm to label a larger dataset of Internet videos.
Acknowledgements
We would like to thank Ilija Radosavovic for helpful discussions and feedback. This work was supported by ONR MURI N00014-21-1-2801. J. W. was supported by the NSF Mathematical Sciences Postdoctoral Fellowship and the UC President’s Postdoctoral Fellowship.

### v3

@@
-Reconstructing Hand-Held Objects in 3D
+Reconstructing Hand-Held Objects in 3D from Images and Videos
 Abstract
-Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from in-the-wild RGB images or videos.
+Objects manipulated by the hand (i.e., manipulanda) are particularly challenging to reconstruct from Internet videos.
 Not only does the hand occlude much of the object, but also the object is often only visible in a small number of image pixels.
@@
 With these insights in mind, we present a scalable paradigm for hand-held object reconstruction that builds on recent breakthroughs in large language/vision models and 3D object datasets.
-Our model, MCC-Hand-Object (MCC-HO), jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs.
-Subsequently, we use GPT-4(V) to retrieve a 3D object model that matches the object in the image and rigidly align the model to the network-inferred geometry; we call this alignment Retrieval-Augmented Reconstruction (RAR).
-Experiments demonstrate that MCC-HO achieves state-of-the-art performance on lab and Internet datasets, and we show how RAR can be used to automatically obtain 3D labels for in-the-wild images of hand-object interactions.
+Given a monocular RGB video, we aim to reconstruct hand-held object geometry in 3D, over time.
+In order to obtain the best performing single frame model, we first present MCC-Hand-Object (MCC-HO), which jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs.
+Subsequently, we prompt a text-to-3D generative model using GPT-4(V) to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR).
+RAR provides unified object geometry across all frames, and the result is rigidly aligned with both the input images and 3D MCC-HO observations in a temporally consistent manner.
+Experiments demonstrate that our approach achieves state-of-the-art performance on lab and Internet image/video datasets.
+We make our code and models available on the project website: https://janehwu.github.io/mcc-ho
+Figure 1: We present a scalable approach to hand-held object reconstruction from monocular RGB images or videos that is guided by object recognition and retrieval.
+Results demonstrate that our method is able to generate realistic object geometry that is also faithful to visual observations and consistent across frames.
+Please see the supplementary materials for evaluation of temporal consistency.
 1 Introduction
-“Man is the measure of all things.”
-Protagoras
-In manipulation, one may philosophize that the hand is the measure of all objects.
-From infancy, humans use their hands to gather information about the physical world [30, 83, 64, 67].
-Through our hands and eyes, we learn to estimate 3D geometry and physical properties as well as affordances that inform the ways in which we manipulate objects [39, 14].
+This paper addresses the problem of recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry.
+Given the tremendous advancements in object recognition and retrieval (e.g., generation), such as GPT-4(V) [51] and text-to-3D models [45, 57, 7, 43], a natural question that arises is whether reconstruction can benefit from leveraging these other branches of computer vision.
+Particularly for the challenging setting of reconstructing hand-held objects, such an approach may be essential to building up available training data to Internet scale.
+See Figure 1 and the supplementary videos.
 Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics.
-While there has been significant progress in reconstructing hands [63, 50, 52] and objects [85, 37, 36] separately, joint reconstruction of hand-object interactions presents unique challenges that cannot currently be solved by general reconstruction approaches.
-Procedural methods such as Structure from Motion (SfM) [72] will struggle because the object is mostly occluded, and thus it is difficult to obtain enough feature points to achieve accurate reconstruction.
-Similarly, Neural Radiance Fields (NeRFs) [44] are best suited for unoccluded multiview reconstruction of static scenes and objects.
-Instead, state-of-the-art approaches in hand-object reconstruction have relied on human assistance for data collection [3, 51, 13].
-While manually-retrieved 3D object models naturally improve reconstruction quality and can provide better initialization, this paradigm is not scalable.
-This is because human-driven retrieval typically includes searching online for a suitable object model, verifying that the model matches the object in the image, and providing an estimate of 3D object pose [3].
-Rigidly aligning the 3D model to an image is also an underconstrained 3D-to-2D problem, and optimization approaches are prone to issues with local minima.
-Figure 1: We present a scalable approach to hand-held object reconstruction that is guided by recognition and retrieval.
-In the first stage, our transformer-based model (MCC-HO) infers hand-object 3D geometry from a single RGB image (Row 2, visualized as meshes).
-In the second stage, a template object model is obtained using Retrieval-Augmented Reconstruction (RAR).
-This template model is then rigidly aligned with the output of MCC-HO using ICP (Row 3).
-In this paper, we present a novel approach to reconstructing hands manipulating common objects from monocular RGB image(s) or video. We draw from three key insights:
-  1. 1.
-    One can recover shape even from single images at test time, if the object being seen has some family resemblance to previously seen 3D objects.
-For children, the ability to see 3D from motion comes first, followed by binocular stereopsis, and then finally they can guess 3D shape even from a single image [92].
-An operationalization of this in computer vision is the MCC system [85] — the model requires 3D information at training time, but at test time a single image is enough.
-  2. 2.
-    In images of hand-held objects, we expect to see hands.
-The hands also provide a strong anchor for 3D location and scale; for instance, the hands provide a convenient measurement tool for many commonplace activities (e.g., cooking, crafts).
-Recent progress in single-image hand reconstruction [52] enables the development of a robust hand-centric approach to object reconstruction.
-  3. 3.
-    Multiple datasets have now been collected which have videos of hand-object interactions with underlying 3D ground truth (see Section 2), but there has not been a model trained on all available data.
-Given an RGB image and estimated 3D hand, we train a transformer-based model, MCC-Hand-Object (MCC-HO), that jointly infers 3D hand and object geometry.
-This geometry is represented as a neural implicit surface composed of occupancy, color, and hand-object segmentation.
-Our model architecture is adapted from an existing object reconstruction model, MCC [85], and thus we make use of a learned object prior by fine-tuning an MCC model that is trained on the CO3Dv2 dataset [59].
-In conjunction with MCC-HO, we propose Retrieval-Augmented Reconstruction (RAR), a method for automatically retrieving object models using large language/vision models to improve network-inferred geometry.
-Specifically, we prompt GPT-4(V) [48] to recognize the hand-held object in the image and provide a detailed text description.
-This description is passed to Genie [42], a text-to-3D generative model, to obtain a 3D object.
-After the object is retrieved, iterative closest point (ICP) [1] or other point cloud fitting techniques can be used to rigidly align the Genie mesh with the estimated geometry from our network.
-See Figure 1.
-To summarize, this paper presents a paradigm for reconstructing hand-held objects that combines the respective advantages of model-free and model-based approaches.
-Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment of the retrieved model with network-inferred geometry.
-In the first stage, MCC-HO jointly estimates hand and object geometry from a single RGB image and 3D hand.
-In the second stage, RAR is used to improve reconstruction quality via recognition and retrieval.
-The hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [42]).
-In the third stage, the retrieved 3D object is rigidly aligned with the network-inferred object geometry.
+First, we aim to solve a particular instance of a classic computer vision problem (along with recognition and segmentation): 3D reconstruction.
+It is also a particularly difficult setting because the hand and object need to be reconstructed from a single RGB image, where the hand is articulated and the object can be small with respect to the image size and partially occluded by the hand.
+For videos, hand-object trajectories need to additionally be temporally consistent.
+Second, in robotics, particularly manipulation, there is a great need for big data.
+Unlike other areas such as natural language processing and computer vision, where big data is already available on the Internet, large-scale trajectories of robots manipulating objects are not readily available.
+There has been a longstanding aspiration to leverage trajectories of human hands interacting with objects [66]; in terms of videos, this data is readily available on the Internet.
+However, the significant challenge is how to extract 3D hand-object trajectories from those videos.
+To this end, our objective is to develop an automated technique for taking raw Internet video and extracting 3D hand-object trajectories in a model-free manner.
+Note that this version of the problem is distinct from model-based approaches that assume a known object shape [55, 100] or use multiview methods to reconstruct the object in 3D in a separate step [13, 31].
+Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment over time of the retrieved model with network-inferred geometry.
+In the first stage, our transformer-based model, MCC-Hand-Object (MCC-HO), estimates hand-held object geometry from each RGB image and 3D hand.
+In the second stage, Retrieval-Augmented Reconstruction (RAR) is used to improve reconstruction quality via recognition and retrieval.
+Specifically, the hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [45]).
+In the third stage, the retrieved 3D object is rigidly aligned with the input image or video using a combination of DINOv2 [52] and inferred 3D cues from MCC-HO.
 Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [4, 2, 41].
-To demonstrate the scalability of our approach, we additionally use RAR to obtain ground truth 3D hands and objects for previously unlabeled images in the 100DOH dataset [66].
+To demonstrate the scalability of our approach, we additionally obtain pseudo ground truth 3D hands and objects for previously unlabeled videos in the 100DOH dataset [71].
+Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs hand-held object trajectories in three stages.
+In the first stage, MCC-HO is used to predict hand and object point clouds.
+In the second stage, a 3D model for the object is obtained using Retrieval-Augmented Reconstruction (RAR).
+In the third stage, the 3D object model is rigidly aligned with DINOv2 [52] image features and network-inferred geometry in a temporally consistent manner.
 2 Related Work
@@
 Hand reconstruction.
-Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [52, 34, 43, 45, 76, 93, 96, 50, 63] or monocular video [28, 82, 88].
+Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [56, 37, 46, 48, 82, 98, 101, 53, 69, 59] or monocular video [30, 87, 92].
 These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [68].
 Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [81, 97].
-Particularly relevant to our work, [61] considers the challenging case of predicting pose when the hand is interacting with objects, which can lead to large occlusions of the hand and/or object.
+We use HaMeR [56] in this work as it has achieved state-of-the-art performance on in-the-wild images.
 Model-based hand-object pose estimation.
 When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [1, 2, 55, 14, 24, 25, 74, 78, 93, 91, 77, 62].
-The estimated hands can then be used to rigidly align the object with images or videos [3, 51] via inverse rendering [41, 58].
-However, image-based optimization losses are prone to issues with local minima, as depth/3D information is ambiguous from a single RGB image.
+The estimated hands can then be used to rigidly align the object with images or videos [2, 55, 100] via inverse rendering [44, 64].
+Due to the specific challenges arising from hand-object interactions, e.g.
+highly occluded geometry, optimization-based approaches are considered the current state-of-the-art in this area.
+RHO [2] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings.
+RHOV [55] extended RHO to video data by adding, among other terms, temporal and optical flow losses.
+Aiming to add more physics-inspired constraints, HandyPriors [100] leverages a differentiable contact module to minimize relative sliding between frames.
+G-HOP [96] guides optimization with a generative hand-object prior.
 Model-free hand-object reconstruction.
 Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions.
-Existing approaches are either tailored to reconstruction from a single image [90, 24, 7, 6, 95, 15, 8], RGB video [27, 91, 19, 53], or RGB-D video [80].
-Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved template 3D model.
+Existing approaches are either tailored to reconstruction from a single image [94, 23, 5, 6, 99, 16, 8, 58], RGB video [28, 95, 21, 96], or RGB-D video [85].
+Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved 3D object model.
 Hand-object datasets.
@@
 Following the release of [71], MOW [2] and RHOV [55] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
-Figure 2: Given an RGB image and estimated 3D hand, our method reconstructs hand-held objects in three stages.
-In the first stage, MCC-HO is used to predict hand and object point clouds.
-In the second stage, a template mesh for the object is obtained using Retrieval-Augmented Reconstruction (RAR).
-In the third stage, the template object mesh is rigidly aligned to the network-inferred geometry using ICP.
 3 Method Overview
-We propose a three-stage approach to reconstructing hand-object interactions that combines feedforward network inference with object model retrieval and subsequent 3D fitting.
+Our approach to reconstructing hand-object interactions from monocular RGB video combines feedforward network inference (MCC-HO) with object model retrieval (RAR) and subsequent rigid alignment.
 See Figure 2.
-Our approach is built on two important insights in the hand-object interaction setting.
-The presence of a human hand provides grounding in terms of both the 3D location and scale of the object.
-While monocular hand reconstruction is an ill-posed problem and can be a limiting factor for hand-conditioned object reconstruction approaches, recent progress has enabled accurate hand estimation by scaling up [52].
+We build on the insight that the presence of a human hand provides grounding in terms of both the 3D location and scale of the object.
 Thus, as in prior work [94, 8], the estimated object pose and shape is conditioned on the estimated 3D hand.
@@
 Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
-In addition, although the hand introduces significant occlusions of the object, the space of possible manipulanda is much smaller than the space of all objects.
-Thus, this work focuses on reconstructing categories of common household objects, similar in spirit to how the COCO dataset [35] was constructed and in alignment with object [60] and hand-object [38] datasets.
-The pervasiveness of such commonplace objects enables deploying large image recognition systems for the task of automatically recognizing the object and retrieving a 3D object model.
-Since household manipulanda are often quite uniform in shape (e.g., a knife, an 8 o.z.
-mug), the accuracy of such a retrieval process is much higher than for generic objects, whether through lookup in an object dataset [11, 10] or text-to-3D generation methods [42] trained on such data.
-4 Hand-Object Geometry Inference
+Given a monocular RGB image or video, our method reconstructs hand-held objects in three stages.
+First, our transformer-based model, MCC-HO, estimates 3D hand-object geometry from single images (Section 4).
+Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [45] through a technique we coin Retrieval-Augmented Reconstruction (Section 5).
+Finally, the 3D object model is rigidly aligned in a temporally consistent manner (Section 6).
+4 Estimating Hand-Object Geometry
 In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [89] for the specific task of hand-object reconstruction.
-Notably, two domain shifts from the original MCC model are applied.
-First, the input depth image is removed and the input RGB-D point cloud is replaced with the 3D hand.
-Second, the model is trained to infer both hand and object geometry, and thus an extra segmentation output is added to the decoder.
-We call our model MCC-Hand-Object (MCC-HO).
+Our model, MCC-Hand-Object (MCC-HO), has to deal with two major challenges.
+First, MCC assumes RGB-D images as input, whereas our model needs to work with RGB input, since our ultimate goal is to reconstruct from Internet videos.
+Second, we have to find a way to cope with the significant domain shift between the training datasets.
+MCC was trained on the CO3D dataset [67], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface.
+For MCC-HO, our training data is a relatively small collection of in-the-wild and lab datasets of hand-object interactions, where the hand-held object is at least partially occluded in the vast majority of images.
 The inputs to MCC-HO are a single RGB image and 3D hand geometry.
-Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM) [29], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image.
-Hand geometry is typically parameterized by the MANO hand model [62], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$.
+Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [32, 65], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image.
+Hand geometry is parameterized by the MANO hand model [68], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$.
 The MANO hand skeleton contains 21 joints.
 At training time, the ground truth 3D hand is used.
-At test time, HaMeR [52] or any other 3D hand estimation methods can be used.
+At test time, HaMeR [56] is used to infer the hand geometry.
 The inferred geometry is represented as a neural implicit function $\rho(x)$ modeling the hand and object jointly, such that for any 3D location $x$, $\rho$ returns the occupancy probability $\sigma(x)$, RGB color $c(x)$, and segmentation label $m(x)$:
@@
 Importantly, $s$ is a constant scaling factor used to ensure the object geometry can be contained by the bounds of this hand-normalized sampling space.
-See Figure 3.
-Figure 3: The input 3D hand is rasterized and per-pixel 3D points are sampled to be passed to the MCC-HO encoder.
 The image $I$ and hand points $P$ are encoded into a single representation $R$, where:
@@
 The decoder architecture is relatively lightweight and also enables dynamic changes to the point sampling resolution and query size at training and/or test time.
-4.2 Implementation Details
-The training loss is a combination of occupancy, color, and segmentation:
+Losses. The training loss is a combination of occupancy, color, and segmentation:
 $\mathcal{L}=\mathcal{L}_{\sigma}+\mathcal{L}_{c}(\sigma_{gt})+\mathcal{L}_{m}$ (5)
@@
 The segmentation loss $\mathcal{L}_{m}$ is 3-way classification loss for each query point corresponding to background (empty space), hand, and object labels.
-Our model was trained on a single Quandro RTX 8000 for 3k epochs with a learning rate of $1e^{-5}$ and batch size of 16.
-We use a cosine schedule and a linear warm-up for the first 5 epochs.
-At each epoch, one random frame from each sequence is selected.
-Network parameters that correspond to the MCC architecture [85] were initialized using a pretrained MCC model that was trained on the CO3Dv2 dataset [60].
-The weights of the additional linear layer in our decoder that outputs hand-object segmentation are initialized via Xavier initialization.
 5 Retrieval-Augmented Reconstruction
-In order to further improve upon the network-inferred object geometry, we use GPT-4(V) [48] to automatically detect and retrieve a 3D object corresponding to the hand-held object in an image.
-We coin this technique Retrieval-Augmented Reconstruction (RAR) in reference to Retrieval-Augmented Generation in natural language processing [32].
-Given an input image (one could also use stacked video frames), we prompt GPT-4(V) to provide a detailed description of the object the hand is holding.
+Like any other deep learning model, how well MCC-HO is expected to perform is related to the amount of training data available, e.g., the size of available hand-object datasets.
+These are currently all quite small, orders of magnitude smaller than the “big data” on which computer vision foundation models have been trained.
+How might we find a way to exploit these other models to our end?
+Specifically, recognition of objects is very well developed [51], and large collections of 3D object models [10, 11] exist and have facilitated advancements in 3D generative modeling [45, 57, 7, 43].
+Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [51] and a text-to-3D generative model [45] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image.
+We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [35].
+Given a representative frame of the video, we prompt GPT-4(V) to provide a detailed description of the object the hand is holding.
 This text caption is passed to Genie, a text-to-3D model from Luma [45], to retrieve realistic 3D object geometry and appearance (Genie is also used in [36]).
-See Figure 4.
-After the object model is obtained, it is rigidly aligned with our network-inferred object geometry using ICP [1].
-In particular, we sample a set of random initial rotations for the retrieved 3D model and select the ICP transformation that results in the lowest alignment error.
-Genie outputs geometry normalized with respect to a unit cube, so the scale of the 3D model is initialized using the bounding volume of the predicted object point cloud.
-Figure 4: Retrieval-Augmented Reconstruction (RAR).
-Given an input image, we prompt GPT-4(V) [48] to recognize and provide a text description of the hand-held object.
-The text description is passed to a text-to-3D model (in our case, Genie [42]) to obtain a 3D object.
+Since Genie outputs geometry normalized to fit inside a unit cube, the scale of the 3D model must be adjusted with respect to the input 3D hand.
+5.1 Rigid Alignment
+After a 3D object model is retrieved, it is rigidly aligned with the input visual data and MCC-HO predictions.
+We begin by considering the single image case.
+First, the object scale is estimated using the predicted MCC-HO object point cloud, which is a reasonable approximation because the prediction is grounded by the input 3D hand.
+If the MCC-HO point cloud is $X\in\mathbb{R}^{N\times 3}$ and a sampled point cloud of the retrieved object model is $Y\in\mathbb{R}^{N\times 3}$, then the scale is defined as $s=\sqrt{\lambda_{X}^{0}}/\sqrt{\lambda_{Y}^{0}}$ where $\lambda_{X}^{0}$ is the largest eigenvalue of $X^{T}X$ and $\lambda_{Y}^{0}$ is similarly defined for $Y$.
+The remaining parameters to be optimized are object rotation $R$ and translation $T$.
+In order to (1) guarantee a globally optimal solution while (2) maintaining fixed compute time, we discretize the space of all possible parameter values to construct finite state spaces $\{\mathcal{R},\mathcal{T}\}$.
+The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [33].
+Translation is defined as an offset from the mean location of the MCC-HO point cloud $\mu(X)$, and the space is discretized by constructing a small voxel grid centered at $\mu(X)$.
+In order to avoid intractably large state space sizes (e.g., $\mathbb{R}^{3}\times\mathbb{R}^{3}$), we first optimize for rotation and then translation.
+For each parameter, we select the state that results in the lowest error using the sum of the following two metrics.
+MCC-HO Chamfer Distance.
+The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric as defined in Equation 8.
+Note that when optimizing for rotation, the object model (centered at the origin) is first rotated and then translated to the mean location of the MCC-HO point cloud.
+DINOv2 PCA Similarity.
+The MCC-HO point cloud gives us a reasonable estimate for object pose, but using this signal alone may not result in poses that align well with the input image (particularly for geometry with symmetries and in-the-wild objects).
+Thus, we additionally use DINOv2 [52] features to determine visual similarities between a transformed object model and the input image.
+For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [64].
+A PCA basis is constructed using the DINOv2 features of the input image, and the first three PCA components are computed for both the rendered features $\mathcal{F}_{j}$ and input image features $\mathcal{F}_{0}$.
+We compute the cosine similarity between the two as
+${\textstyle E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)}$ (6)
+where $\mathcal{F}_{j}$ and $\mathcal{F}_{0}$ are both masked by the input image object silhouette.
+When ground truth object masks are not available, we use SAM 2 [65].
+See Figure 3 for rotation results using example images from 100DOH [71].
+Figure 3: We compute a PCA basis of DINOv2 features using all frames masked by the object silhouettes (one frame and its first three PCA components are shown on the left side of each example).
+The first three components of this basis are used to determine the maximum likelihood Genie object rotation for each frame (shown on the right side).
+6 From Images to Videos
+In the previous section, we showed how given a single image, we select the rotation and translation of the 3D object model that minimizes error metrics based on CD and DINOv2 similarity.
+Now, if we have as input a video stream, we can further exploit the additional cue of temporal smoothness—we expect adjacent frames to have rotation and translation states that are close to each other.
+Given a sequence of video frames with timesteps $t$, we jointly optimize for per-frame rotations $R(t)$ and subsequently per-frame translations $T(t)$.
+The typical approach for temporal alignment is to use gradient-based optimization  [55, 95], which is prone to issues with local minima and has unbounded computation time.
+Aiming for robustness and efficiency, we propose using the Viterbi algorithm [80, 15, 29] to solve for a global solution to rigid object alignment across all frames of a video sequence.
+The Viterbi algorithm is a dynamic programming approach to obtaining the maximum a posteriori probability of the most likely sequence of hidden states $Q=q_{1},q_{2},\ldots,q_{N}$ that results in a sequence of observed events $O=o_{1},o_{2},\ldots,o_{N}$.
+The hidden states $q_{t}$ can take on values from a finite set of parameters $\mathcal{Q}$ (where in our case, $\mathcal{Q}=\mathcal{R}$ or $\mathcal{Q}=\mathcal{T}$).
+Following the single image case, the observations $o_{t}$ are a combination of the MCC-HO inferred object point clouds and the DINOv2 image features.
+Each cell of the Viterbi trellis, $v_{t}(j)$, is the probability of ending up in state $j$ after seeing the first $t$ observations and passing through the most probable state sequence $q_{1},\ldots,q_{t-1}$.
+Given the probability of being in every state at time $t-1$, the Viterbi probabilities are computed via:
+$v_{t}(j)=\max_{i=1}^{|\mathcal{Q}|}\hskip 2.84526ptv_{t-1}(i)a_{ij}b_{j}(o_{t}),$ (7)
+where $a_{ij}$ is the transition probability from the previous state $q_{i}$ to the current state $q_{j}$, and $b_{j}(o_{t})$ is the state observation likelihood given the current state $j$.
+In particular, $a_{ij}$ is Rodrigues’ rotation error when optimizing for rotations and Euclidean distance when optimizing for translations.
+The emission cost $b_{j}(o_{t})$ is the sum of the two metrics defined in the single image discussion, e.g., CD with respect to the MCC-HO point clouds and DINOv2 PCA similarity with respect to the input images.
 7 Experiments
@@
 Section 7.2 defines the evaluation metrics used for comparisons.
-Section 6.3 quantifies the advantages of using MCC-HO, e.g., a hand-centric approach, in comparison to the original MCC model [85].
-We also quantitatively (Section 6.4) and qualitatively (Section 6.5) evaluate our model against existing work in model-free, monocular hand-object reconstruction [24, 90] and find that MCC-HO achieves state-of-the-art performance on all three datasets used [5, 3, 38].
-In particular, the methods most similar to ours are HO [24] and IHOI [90].
-The quality of the MCC-HO predicted geometry is also shown to be sufficient for accurate rigid alignment of template objects using ICP.
-Also in Section 6.5.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach on in-the-wild Internet images to demonstrate the feasibility and scalability of our automated model retrieval and fitting.
-Specifically, we apply MCC-HO and RAR to unlabeled images from the 100DOH dataset [66].
+We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [4, 2, 41].
+Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [71] to demonstrate the feasibility and scalability of our automated model retrieval and fitting.
+Section 7.5 includes ablation studies for MCC-HO, RAR, and each component of our Viterbi temporal alignment algorithm.
+DexYCB MOW HOI4D
+F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
+HO [23] 0.24 0.48 4.76 0.03 0.06 49.8 0.28 0.51 3.86
+IHOI [94] - - - 0.13 0.24 23.1 0.42 0.70 2.7
+HORSE [58] 0.20 0.41 7.16 0.08 0.20 27.2 0.21 0.41 7.68
+MCC-HO (Ours) 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
+Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D. Chamfer Distance (cm2) and F-score (5mm, 10mm) are reported.
+Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
+DiffHOI [89] 0.62 0.91 0.8
+G-HOP [96] 0.76 0.97 0.4
+MCC-HO + RAR (Ours) 0.91 0.98 0.28
+Table 2: Quantitative evaluation of video-based hand-held object reconstruction methods on the HOI4D test dataset. Mean metrics are reported for consistency with prior work.
 7.1 Datasets
 We experiment on four different datasets for hand-object reconstruction.
-HOI4D: The HOI4D dataset [38] consists of 4,000 lab videos of hands interacting with specified categories of common objects.
+HOI4D.
+The HOI4D dataset [41] consists of 4,000 lab videos of hands interacting with specified categories of common objects.
 We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets.
@@
 For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
-DexYCB: The DexYCB dataset [5] contains 1,000 lab videos of hands interacting with 21 YCB objects.
-We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences (consistent with prior work [7, 6]).
-MOW/RHOV: MOW [3] is a dataset of 512 images from 100 Days of Hands [66] labeled with 3D hands and objects.
+DexYCB. The DexYCB dataset [4] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
+MOW/RHOV.
+MOW [2] is a dataset of 512 images from 100 Days of Hands [71] labeled with 3D hands and objects.
 We use the same splits as [94] with 350 training images and 92 test images.
@@
 In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
-100DOH: The 100 Days of Hands dataset [66] contains 131 days of footage collected from Internet videos showing hands interacting with objects.
+100DOH.
+The 100 Days of Hands dataset [71] contains 131 days of footage collected from Internet videos showing hands interacting with objects.
 The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
@@
 7.2 Evaluation Metrics
-Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics, Chamfer Distance (CD) and F-score reported at 5mm and 10mm.
-Both metrics are computed on point sets, and thus the point clouds inferred by MCC-HO are directly used.
+Since the focus of our approach is on the object manipulated by the hand, we quantitatively evaluate object reconstruction quality using standard 3D metrics: CD (cm2) and F-score reported at 5mm and 10mm.
+Both metrics are computed on point sets.
 The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [69] and HOI4D optimizes hand pose using 2D keypoint annotations.
@@
 In order to account for unknown ground truth camera parameters, ICP with scaling [64] is used to align all predicted geometry with the ground truth mesh.
-Median metrics are reported for each test dataset.
-Chamfer Distance: The Chamfer Distance metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
+Median metrics are reported unless otherwise noted.
+Chamfer Distance. The CD metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
 $CD(A,B)=\frac{1}{N}\sum_{a\in A}\min_{b\in B}||a-b||^{2}_{2}+\frac{1}{N}\sum_{b\in B}\min_{a\in A}||a-b||^{2}_{2},$ (8)
-[…]
-F-score: For 3D evaluation of predicted points, the F-score at a given threshold $t$ is defined as:
-$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{\text{Precision}(X_{pd},X_{gt},t)\cdot\text{Recall}(X_{pd},X_{gt},t)}{\text{Precision}(X_{pd},X_{gt},t)+\text{Recall}(X_{pd},X_{gt},t)}.$ (7)
-6.3 MCC Baseline
-Because our model is built on top of MCC [85], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction.
-For MCC, we use the released model that was trained on the CO3Dv2 dataset [60] and apply an existing monocular depth estimator [57] to obtain depth maps (as MCC takes an RGB-D image as input).
-Table 1 compares 3D metrics between MCC and MCC-HO computed on the MOW test set.
-We find that fine-tuning MCC for the specific task of hand-held object reconstruction improves performance on test images unseen by both models.
-Model Parameters F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
-MCC [85] 0.10 0.21 31.6
-MCC-HO 0.15 0.31 15.2
-[…]
-DexYCB MOW HOI4D
-F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
-HO [24] 0.24 0.48 4.76 0.03 0.06 49.8 0.28 0.51 3.86
-IHOI [90] - - - 0.13 0.24 23.1 0.42 0.70 2.7
-MCC-HO 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
-[…]
+where the point sets should have the same size. This equation is consistent with prior work [5, 6, 94, 95]. We report CD in cm2 as in[5, 6].
+F-score. The F-score at a given threshold $t$ is defined as a function of precision $P$ and recall $R$:
+$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{P(X_{pd},X_{gt},t)\cdot R(X_{pd},X_{gt},t)}{P(X_{pd},X_{gt},t)+R(X_{pd},X_{gt},t)}.$ (9)
 7.3 Quantitative Evaluation
-MCC-HO is quantitatively evaluated on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [24, 90].
+We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [23, 94, 58].
 See Table 1.
 In order to sample points more densely in hand-object regions, the dimensions and granularity of the voxel grid used to query MCC-HO are determined from an initial coarse prediction using a default voxel grid.
-For DexYCB, we use every 5th frame of the “s0” test split, and for HOI4D, we use all frames from the 12 test sequences that contain a ground truth 3D hand label.
-The results for IHOI [90] and HO [24] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [91] papers.
+For DexYCB and HOI4D, we use every 5th frame of each test video sequence.
+The results for IHOI [94] and HO [23] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [95] papers; otherwise, we used available public repositories of prior work.
 Because AlignSDF [5] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics.
@@
 In addition, the pretrained MCC model provides a strong object prior for initialization.
+We also compare to existing video-based approaches for RGB hand-held object reconstruction in Table 2, noting that the ground truth object is used for our RAR rigid alignment (a reasonable comparison as DiffHOI [95] and GHOP [96] train category-specific diffusion models using the ground truth HOI4D 3D objects).
 7.4 Qualitative Evaluation
-We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well template objects can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks).
-Figure 5 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud using ICP.
+We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well retrieved object models using RAR can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks).
+Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud.
 Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera.
 Because the object geometry is conditioned on the 3D hand, our approach tends to predict reasonable hand-object contact.
-Additionally, the object in contact with the hand is correctly identified by MCC-HO despite being provided a hand-object bounding box and no segmentation maps.
+Additionally, the object in contact with the hand is correctly identified by MCC-HO without input segmentation maps.
 Figure 4: MCC-HO results for MOW test examples.
-The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud using ICP (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
-Figure 6: Our approach applied to test images from MOW (Rows 1-3), DexYCB (Row 4), and HOI4D (Row 5).
-Each example includes the input image (left), rigid alignment of the ground truth mesh with the MCC-HO predicted point cloud using ICP (middle), and an alternative view (right).
+The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
 Figure 5: Our approach applied to images in the 100 Days of Hands dataset [71].
 For each image, we estimate 3D hands using HaMeR [56], inference object geometry using MCC-HO, retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie), and rigidly align the 3D model with the inferred point cloud from MCC-HO.
-Figure 6 shows additional visualizations of applying rigid alignment of the template model to MCC-HO results.
-We include held-out test examples from all three datasets used for training: MOW, DexYCB, and HOI4D.
-Because MCC-HO is fine-tuned from a general object reconstruction transformer model (MCC), it is robust to unseen object instances at test time.
-For instance, objects in the MOW test images were never seen by MCC-HO during training.
-From a practical standpoint, we find 3D-to-3D fitting of template objects with the MCC-HO predicted geometry (e.g., via ICP) to be significantly more efficient than prior 3D-to-2D or per-example optimization approaches [3, 51, 91].
-Moreover, optimizing 3D pose with respect to 2D images tends to get stuck in local minima.
-As Figure 6 demonstrates, this 3D fitting results in physically plausible hand-object interactions that when projected align with the image data.
-6.5.1 Retrieval-Augmented Reconstruction.
-The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images.
-Specifically, we take 9 images from 100DOH [66] that do not have 3D ground truth (e.g., not in MOW) and apply MCC-HO combined with RAR to reconstruct hand-object 3D geometry.
-See Figure 7.
+7.4.1 Retrieval-Augmented Reconstruction
+The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images and videos.
+Figure 5 shows our results for 9 images from 100DOH [71] that do not have 3D ground truth (e.g., not in MOW).
 Estimated 3D hands for each image are obtained using HaMeR [56].
 At test time, our method does not require an input hand-object segmentation mask [2, 55], depth image [89], or hand joint locations [94, 8].
-Rather, we simply run feed-forward inference using MCC-HO, retrieve a 3D object mesh using RAR, and run ICP to align this mesh with the inferred point cloud.
-As more training data is collected through RAR and human-in-the-loop verification, we expect MCC-HO to naturally improve as well (similar in spirit to the data collection process used in Segment Anything [29]).
-6.6 Limitations
-While our 3D-to-3D rigid alignment approach using ICP is efficient, some misalignment may occur when rendering the reconstruction from the input view camera.
-This is because we do not use image-based cues for alignment, but note that this can be easily fixed via inverse rendering as in [3, 51].
-In particular, the 6DOF object pose can be fine-tuned by minimizing the silhouette loss and hand-object contact/collision losses defined in [3].
-See Figure 8.
-Regardless, our reconstruction provides a more automatic/robust initialization of the 3D object model than manual or random initialization, as was done in prior work [3, 51].
-In addition, because MCC-HO is conditioned on an input 3D hand, the accuracy of the predicted hand pose naturally affects object reconstruction performance.
-Figure 8: The input image is not used when running ICP, which can sometimes result in image misalignment errors (Row 1).
-A postprocess can be applied to resolve silhouette misalignment and any hand-object contact issues (Row 2).
+Please see the supplementary materials for additional results.
+7.5 Ablation Studies
+MCC Baseline.
+Because our model is built on top of MCC [89], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction.
+For MCC, we use the released model that was trained on the CO3Dv2 dataset [67] and apply an existing monocular depth estimator [63] to obtain depth maps (as MCC takes an RGB-D image as input).
+Table 3 compares 3D metrics between MCC and MCC-HO computed on the MOW test set.
+We find that using MCC-HO, which was trained on the specific task of hand-held object reconstruction, leads to performance improvements on test images.
+Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
+MCC [89] 0.10 0.21 31.6
+MCC-HO 0.15 0.31 15.2
+Table 3: Analysis of how hand-held object reconstruction benefits from conditioning on the hand using the MOW test dataset.
+RAR Success Rate.
+In order to quantify the success rate of RAR (e.g., whether a generated 3D object matches the input image), we use GPT-4(V) and Genie to automatically generate objects for all 512 images in the MOW dataset.
+Genie is able to generate an object 82% of the time, and otherwise returned an error.
+The results are included in the supplementary materials.
+Video Alignment Components.
+We aim to understand the contributions of each cost term in our Viterbi algorithm, e.g., MCC-HO and DINOv2.
+See Table 4.
+In order to best assess the quality of rigid alignment, the ICP postprocess described in Section 7.2 is not applied.
+The results demonstrate that both terms contribute to the efficacy of our Viterbi algorithm; in practice, we observe that using both 2D and 3D cues is particularly important for objects that lie outside the training distribution of MCC-HO (DexYCB consists of commonplace YCB objects).
+Viterbi Cost Terms F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
+MCC-HO 0.26 0.52 4.18
+DINOv2 0.23 0.45 6.38
+MCC-HO + DINOv2 0.27 0.52 4.10
+Table 4: Analysis of how each of our Viterbi cost terms contributes to the final temporal alignment on the DexYCB test dataset.
 8 Conclusion
@@
 We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets.
-These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH images [66].
-The combination of MCC-HO and RAR can enable the construction of large hand-object datasets at-scale; for future work, we plan to expand on this paradigm to label a larger dataset of Internet videos.
+These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [71].
+Limitations.
+As 3D generative models improve, RAR will naturally be able to retrieve more accurate and realistic geometry.
+Our approach also relies on the ability to obtain accurate hand and object masks which are tracked across frames.
+We expect that both MCC-HO and our rigid alignment algorithm will reap the benefits of any future image segmentation/tracking improvements.
 Acknowledgements

### v4

@@
 In order to obtain the best performing single frame model, we first present MCC-Hand-Object (MCC-HO), which jointly reconstructs hand and object geometry given a single RGB image and inferred 3D hand as inputs.
-Subsequently, we prompt a text-to-3D generative model using GPT-4(V) to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR).
+Subsequently, we prompt a text-to-3D generative model using a VLM to retrieve a 3D object model that matches the object in the image(s); we call this alignment Retrieval-Augmented Reconstruction (RAR).
 RAR provides unified object geometry across all frames, and the result is rigidly aligned with both the input images and 3D MCC-HO observations in a temporally consistent manner.
@@
 1 Introduction
-This paper addresses the problem of recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry.
-Given the tremendous advancements in object recognition and retrieval (e.g., generation), such as GPT-4(V) [51] and text-to-3D models [45, 57, 7, 43], a natural question that arises is whether reconstruction can benefit from leveraging these other branches of computer vision.
-Particularly for the challenging setting of reconstructing hand-held objects, such an approach may be essential to building up available training data to Internet scale.
+Recovering 3D hand-object interactions from visual data is an important problem in both computer vision and robotics.
+While there has been significant progress in reconstructing hands [pavlakos2023reconstructing, potamias2025wilor, ye2025predicting] and objects [liu2023zero, liu2024one, long2024wonder3d, xu2024instantmesh] separately, joint reconstruction of hand-object interactions presents unique challenges that cannot currently be solved by general reconstruction approaches.
+Existing image-to-3D models [liu2023zero, liu2024one, long2024wonder3d, xu2024instantmesh] are trained on data where the object is usually fully visible in the image; thus, these models are not trained to infer geometry primarily from hand/scene context.
+Procedural methods such as Structure from Motion (SfM) [tomasi1992shape] will also struggle because the object is mostly occluded, and thus it is difficult to obtain enough feature points to achieve accurate reconstruction.
+Instead, state-of-the-art approaches in hand-object reconstruction have relied on human assistance for data collection, whether it be lab datasets [chao2021dexycb, liu2022hoi4d, fan2023arctic] or Internet videos [cao2021reconstructing, patel2022learning].
+Manually scanned or retrieved object models naturally improve reconstruction quality and can provide better initialization, but this paradigm is not scalable.
+In this paper, we present a novel approach to recovering hand-object interactions in 3D from monocular RGB images and videos, focusing in particular on reconstructing hand-held object geometry.
 See Figure 1 and the supplementary videos.
-Recovering hand-object interactions in 3D is an important problem in both computer vision and robotics.
-First, we aim to solve a particular instance of a classic computer vision problem (along with recognition and segmentation): 3D reconstruction.
-It is also a particularly difficult setting because the hand and object need to be reconstructed from a single RGB image, where the hand is articulated and the object can be small with respect to the image size and partially occluded by the hand.
-For videos, hand-object trajectories need to additionally be temporally consistent.
-Second, in robotics, particularly manipulation, there is a great need for big data.
-Unlike other areas such as natural language processing and computer vision, where big data is already available on the Internet, large-scale trajectories of robots manipulating objects are not readily available.
-There has been a longstanding aspiration to leverage trajectories of human hands interacting with objects [66]; in terms of videos, this data is readily available on the Internet.
-However, the significant challenge is how to extract 3D hand-object trajectories from those videos.
-To this end, our objective is to develop an automated technique for taking raw Internet video and extracting 3D hand-object trajectories in a model-free manner.
-Note that this version of the problem is distinct from model-based approaches that assume a known object shape [55, 100] or use multiview methods to reconstruct the object in 3D in a separate step [13, 31].
-Our method consists of three stages: (1) learning-based hand-object reconstruction, (2) object model retrieval, and (3) rigid alignment over time of the retrieved model with network-inferred geometry.
-In the first stage, our transformer-based model, MCC-Hand-Object (MCC-HO), estimates hand-held object geometry from each RGB image and 3D hand.
-In the second stage, Retrieval-Augmented Reconstruction (RAR) is used to improve reconstruction quality via recognition and retrieval.
-Specifically, the hand-held object is recognized by GPT-4(V) and a 3D object is retrieved using a text-to-3D generative model (e.g., Genie [45]).
-In the third stage, the retrieved 3D object is rigidly aligned with the input image or video using a combination of DINOv2 [52] and inferred 3D cues from MCC-HO.
-Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [4, 2, 41].
-To demonstrate the scalability of our approach, we additionally obtain pseudo ground truth 3D hands and objects for previously unlabeled videos in the 100DOH dataset [71].
-Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs hand-held object trajectories in three stages.
-In the first stage, MCC-HO is used to predict hand and object point clouds.
-In the second stage, a 3D model for the object is obtained using Retrieval-Augmented Reconstruction (RAR).
-In the third stage, the 3D object model is rigidly aligned with DINOv2 [52] image features and network-inferred geometry in a temporally consistent manner.
+Given an RGB image and estimated 3D hand, we train a transformer-based model, MCC-Hand-Object (MCC-HO), that jointly infers 3D hand and object geometry.
+This geometry is represented as a neural implicit surface composed of occupancy, color, and hand-object segmentation.
+By adapting our model architecture from an existing object reconstruction model, MCC [wu2023multiview], we make use of a learned object prior by fine-tuning an MCC model that is pretrained on the CO3Dv2 dataset [reizenstein2021common].
+Experiments show that our model outperforms existing methods for hand-held object reconstruction on available datasets containing 3D labels, including DexYCB, MOW, and HOI4D [chao2021dexycb, cao2021reconstructing, liu2022hoi4d] (Table 1, Figures 4 and 5).
+Given the limited size of existing 3D hand-object datasets, we additionally leverage vision-language models [gpt4v, comanici2025gemini] and 3D generative models [luma, poole2022dreamfusion, chen2024text, long2024wonder3d] in order to improve upon a domain-specific approach.
+In conjunction with MCC-HO, we propose Retrieval-Augmented Reconstruction (RAR), a method for automatically retrieving object models using large language/vision models to improve network-inferred geometry.
+Specifically, we prompt a VLM [gpt4v] to recognize the hand-held object in the image and provide a detailed text description.
+This description is passed to a text-to-3D generative model [luma] to obtain a 3D object, which is then rigidly aligned with the input image or video using a combination of DINOv2 [oquab2023dinov2] and inferred 3D cues from MCC-HO.
+We demonstrate that RAR leads to quantitative improvements in 3D reconstruction (Tables 2 and 5), as well as more realistic object geometries compared to existing image-to-3D methods (Figures 5 and 6).
+Note that while RAR works well on average, it will fail sometimes if no close enough 3D match is found by the generative model.
+But the good news is that it does not need to work well on every video for the results of this framework to be useful as a “data engine” for robotics.
+There is a great deal of interest in getting hand-object interaction trajectories for imitation learning by robots [singh2024hand] and typically for this data is collected in lab settings [banerjee2024introducing, fu2025gigahands, hoque2025egodex] which does not result in sufficient diversity.
+With our paradigm, we can use a quick “thumbs-up, thumbs-down” check from a human observer to screen for all videos where the human-object interactions are reconstructed good enough.
+We have applied this approach with great success to Internet videos from the 100DOH dataset [shan2020understanding]; see the supplementary materials.
+While these videos will be of slightly lower quality than those in laboratory collected datasets, they have the advantage of scale and diversity and thus can constitute the pre-training phase of a robot data pipeline with fine-tuning provided by the laboratory datasets.
+Figure 2: Given an RGB video and estimated 3D hands, our method reconstructs 3D hand-held object trajectories.
+First, MCC-HO is used to predict hand and object point clouds for each frame (Section 4).
+Then, a single 3D model for the object is obtained using Retrieval-Augmented Reconstruction (Section 5).
+The 3D object model is rigidly aligned with DINOv2 [oquab2023dinov2] image features and network-inferred geometry in a temporally consistent manner via our Viterbi algorithm (Section 6).
 2 Related Work
@@
 Reconstructing objects from images is a longstanding problem in computer vision, and the bulk of research in this area has focused on scenarios where the object is fully or almost fully unoccluded in the input view(s).
-Structure from motion (SfM) [22, 70, 72, 76] and SLAM [3, 50, 73] remain robust solutions particularly for multiview reconstruction and camera parameter estimation.
-Learning-based object reconstruction methods typically represent 3D geometry as point clouds [60, 61], meshes [17, 86, 88], voxel grids [90, 83], or implicit surfaces [54, 47, 49, 75, 84, 89].
-Following the release of Objaverse [10] and Objaverse-XL [11], a number of works have used this large-scale 3D object dataset to train monocular object reconstruction models [40, 39, 42, 27].
-The CO3Dv2 dataset [67] contains objects from 50 MS-COCO categories [38] and has also been used to train monocular object reconstruction models [89].
+Structure from motion (SfM) [hartley2003multiple, scharstein2002taxonomy, szeliski2022computer, tomasi1992shape] and SLAM [castellanos1999spmap, mur2015orb, taketomi2017visual] remain robust solutions particularly for multiview reconstruction and camera parameter estimation.
+Learning-based object reconstruction methods typically represent 3D geometry as point clouds [qi2017pointnet, qi2017pointnet++], meshes [gkioxari2019mesh, wen2019pixel2mesh++, worchel2022multi], voxel grids [xie2019pix2vox, wang2021multi], or implicit surfaces [park2019deepsdf, mildenhall2021nerf, muller2022instant, tewari2020state, wang2021neus, wu2023multiview].
+Following the release of Objaverse [deitke2023objaverse] and Objaverse-XL [deitke2024objaverse], a number of works have used this large-scale 3D object dataset to train monocular object reconstruction models [liu2023zero, liu2024one, long2023wonder3d, hong2023lrm].
+The CO3Dv2 dataset [reizenstein21co3d] contains objects from 50 MS-COCO categories [lin2014microsoft] and has also been used to train monocular object reconstruction models [wu2023multiview].
 In practice, we find that off-the-shelf object reconstruction methods do not perform well on images of hand-object interactions due to occlusion from the hand, larger object pose diversity, etc.
 Hand reconstruction.
-Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [56, 37, 46, 48, 82, 98, 101, 53, 69, 59] or monocular video [30, 87, 92].
-These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [68].
-Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [81, 97].
-We use HaMeR [56] in this work as it has achieved state-of-the-art performance on in-the-wild images.
+Recently, there has been a growing body of literature on estimating hand pose from a monocular RGB image [pavlakos2023reconstructing, li2022interacting, meng20223d, moon2023bringing, wang2023memahand, yu2023acr, zuo2023reconstructing, park2022handoccnet, rong2020frankmocap, prakash20243d] or monocular video [Karunratanakul_2023_CVPR, Wen_2023_CVPR, Xu_2023_CVPR].
+These approaches either estimate skeletal joint locations of the hands or use parametric models such as MANO [romero2022embodied].
+Since humans have two hands, other methods aim to reconstruct the left and right hands jointly [Wang_2023_CVPR, Yu_2023_CVPR].
+We use HaMeR [pavlakos2023reconstructing] in this work as it has achieved state-of-the-art performance on in-the-wild images.
 Model-based hand-object pose estimation.
-When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [1, 2, 55, 14, 24, 25, 74, 78, 93, 91, 77, 62].
-The estimated hands can then be used to rigidly align the object with images or videos [2, 55, 100] via inverse rendering [44, 64].
+When 3D object models are available, the problem of recovering hand-object interactions can be reduced to estimating the pose and shape of the hand and the rigid pose of the object [bhatnagar2022behave, cao2021reconstructing, patel2022learning, fan2023arctic, hasson2020leveraging, hasson2021towards, tekin2019h+, tzionas2016capturing, yang2022artiboost, xie2022chore, tse2022collaborative, qu2023novel].
+The estimated hands can then be used to rigidly align the object with images or videos [cao2021reconstructing, patel2022learning, zhang2023handypriors] via inverse rendering [loper2014opendr, ravi2020accelerating].
 Due to the specific challenges arising from hand-object interactions, e.g.
 highly occluded geometry, optimization-based approaches are considered the current state-of-the-art in this area.
-RHO [2] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings.
-RHOV [55] extended RHO to video data by adding, among other terms, temporal and optical flow losses.
-Aiming to add more physics-inspired constraints, HandyPriors [100] leverages a differentiable contact module to minimize relative sliding between frames.
-G-HOP [96] guides optimization with a generative hand-object prior.
+RHO [cao2021reconstructing] studied the problem of recovering hand-object pose from in-the-wild images, whereas prior work was centered around datasets captured in lab settings.
+RHOV [patel2022learning] extended RHO to video data by adding, among other terms, temporal and optical flow losses.
+Aiming to add more physics-inspired constraints, HandyPriors [zhang2023handypriors] leverages a differentiable contact module to minimize relative sliding between frames.
+G-HOP [ye2024g] guides optimization with a generative hand-object prior.
 Model-free hand-object reconstruction.
 Reconstructing hand-held objects without corresponding 3D models poses additional challenges compared to general object reconstruction due to frequent occlusions by the hand and significantly less publicly available 3D data of hand-object interactions.
-Existing approaches are either tailored to reconstruction from a single image [94, 23, 5, 6, 99, 16, 8, 58], RGB video [28, 95, 21, 96], or RGB-D video [85].
+Existing approaches are either tailored to reconstruction from a single image [ye2022s, hasson2019learning, chen2022alignsdf, chen2023gsdf, zhang2023moho, fu2023hacd, choi2023handnerf, Prakash2024HOI], RGB video [huang2022reconstructing, ye2023diffusion, hampali2023hand, ye2024g], or RGB-D video [wen2023bundlesdf].
 Our approach combines the respective benefits of model-based and model-free reconstruction; we first reconstruct 3D objects from a single image, and the output is used to guide the registration of a retrieved 3D object model.
 Hand-object datasets.
-Datasets containing hand-object interactions are either collected in lab environments [1, 20, 4, 34, 41], through large-scale capture [18, 19], or from Internet videos [9, 71, 2].
-A number of hand-held object reconstruction models [5, 6, 8] were trained on either the HO-3D [20] or DexYCB [4] datasets.
+Datasets containing hand-object interactions are either collected in lab environments [bhatnagar2022behave, hampali2020honnotate, chao2021dexycb, kwon2021h2o, liu2022hoi4d], through large-scale capture [grauman2022ego4d, grauman2023ego], or from Internet videos [damen2018scaling, shan2020understanding, cao2021reconstructing].
+A number of hand-held object reconstruction models [chen2022alignsdf, chen2023gsdf, choi2023handnerf] were trained on either the HO-3D [hampali2020honnotate] or DexYCB [chao2021dexycb] datasets.
 Since the combined number of object instances in these two datasets is 31, it is difficult for such approaches to generalize to unseen objects.
-HOI4D [41] is currently the largest available dataset with 3D hand and object labels (also used in [95]).
+HOI4D [liu2022hoi4d] is currently the largest available dataset with 3D hand and object labels (also used in [ye2023diffusion]).
 There is also a comparably larger amount of data containing videos of hand-object interactions without 3D labels.
-Epic Kitchens [9] contains 55 hours of egocentric video of people in their personal kitchens.
-Ego4D [18] and Ego-Exo4D [19] are also egocentric datasets, and 100 Days of Hands (100DOH) [71] contains 131 days of footage collected from Internet videos.
-Following the release of [71], MOW [2] and RHOV [55] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
+Epic Kitchens [damen2018scaling] contains 55 hours of egocentric video of people in their personal kitchens.
+Ego4D [grauman2022ego4d] and Ego-Exo4D [grauman2023ego] are also egocentric datasets, and 100 Days of Hands (100DOH) [shan2020understanding] contains 131 days of footage collected from Internet videos.
+Following the release of [shan2020understanding], MOW [cao2021reconstructing] and RHOV [patel2022learning] labeled a subset of images and video clips from 100DOH, respectively, with 3D hands and objects.
 3 Method Overview
@@
 See Figure 2.
-We build on the insight that the presence of a human hand provides grounding in terms of both the 3D location and scale of the object.
-Thus, as in prior work [94, 8], the estimated object pose and shape is conditioned on the estimated 3D hand.
-Applying a hand-centric approach has several advantages.
-Since monocular image/video reconstruction is particularly ill-posed with respect to depth and scale, 3D hands can be used to disambiguate the location and size of the object.
-Furthermore, the location of the 3D hand provides a coordinate system for estimating object geometry; in the case of rotating, symmetric objects, the hand can disambiguate the pose of the object.
-Given a monocular RGB image or video, our method reconstructs hand-held objects in three stages.
 First, our transformer-based model, MCC-HO, estimates 3D hand-object geometry from single images (Section 4).
-Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [45] through a technique we coin Retrieval-Augmented Reconstruction (Section 5).
+Simultaneously, a 3D object model is automatically synthesized by a text-to-3D generative model [luma] through a technique we coin Retrieval-Augmented Reconstruction (Section 5).
 Finally, the 3D object model is rigidly aligned in a temporally consistent manner (Section 6).
 4 Estimating Hand-Object Geometry
-In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [89] for the specific task of hand-object reconstruction.
+In the first stage, hand and object geometry are jointly inferred by training a model adapted from MCC [wu2023multiview] for the specific task of hand-object reconstruction.
 Our model, MCC-Hand-Object (MCC-HO), has to deal with two major challenges.
@@
 Second, we have to find a way to cope with the significant domain shift between the training datasets.
-MCC was trained on the CO3D dataset [67], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface.
+MCC was trained on the CO3D dataset [reizenstein21co3d], which consists of clean, 360-degree videos where each object is unoccluded and placed on a solid surface.
 For MCC-HO, our training data is a relatively small collection of in-the-wild and lab datasets of hand-object interactions, where the hand-held object is at least partially occluded in the vast majority of images.
 The inputs to MCC-HO are a single RGB image and 3D hand geometry.
-Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [32, 65], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image.
-Hand geometry is parameterized by the MANO hand model [68], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$.
+Hand and object segmentation masks can be obtained either from ground truth labels or Segment Anything (SAM/SAM2) [kirillov2023segment, ravi2024sam], and the combined hand-object mask determines the bounding box that is used to crop/resize the input image.
+Hand geometry is parameterized by the MANO hand model [romero2022embodied], which is composed of pose parameters $\theta\in\mathbb{R}^{48}$ and shape parameters $\beta\in\mathbb{R}^{10}$.
 The MANO hand skeleton contains 21 joints.
 At training time, the ground truth 3D hand is used.
-At test time, HaMeR [56] is used to infer the hand geometry.
+At test time, HaMeR [pavlakos2023reconstructing] is used to infer the hand geometry.
 The inferred geometry is represented as a neural implicit function $\rho(x)$ modeling the hand and object jointly, such that for any 3D location $x$, $\rho$ returns the occupancy probability $\sigma(x)$, RGB color $c(x)$, and segmentation label $m(x)$:
@@
 4.1 Architecture
-MCC-HO has an encoder-decoder architecture that is composed of transformers [79] and outputs an implicit representation of hand-object geometry.
+MCC-HO has an encoder-decoder architecture that is composed of transformers [vaswani2017attention] and outputs an implicit representation of hand-object geometry.
 The input images and 3D hand are passed to the encoder to compute a feature map that is a function of both image and hand embeddings.
@@
 The encoder takes as input a single RGB image $I\in\mathbb{R}^{H\times W\times 3}$ and a posed MANO hand mesh defined in camera world space with vertices $v_{h}\in\mathbb{R}^{778\times 3}$ and faces $f_{h}\in\mathbb{R}^{1538\times 3}$.
-The camera intrinsics are either known or consistent with the estimated 3D hand (e.g., the camera parameters used during HaMeR inference [56]).
+The camera intrinsics are either known or consistent with the estimated 3D hand (e.g., the camera parameters used during HaMeR inference [pavlakos2023reconstructing]).
 In order to jointly encode the image and hand geometry, the hand mesh is rasterized and per-pixel 3D points for the hand are sampled.
@@
 $R\coloneqq f(E^{RGB}(I),E^{XYZ}(P))\in\mathbb{R}^{N^{enc}\times C},$ (3)
-and $E^{RGB}$ and $E^{XYZ}$ are identical to the transformers proposed in MCC [89], though we define $P$ differently, i.e., without an input depth image.
-The image transformer $E^{RGB}$ has a Vision Transformer (ViT) architecture [12] with $16\times 16$ patch embeddings.
+and $E^{RGB}$ and $E^{XYZ}$ are identical to the transformers proposed in MCC [wu2023multiview], though we define $P$ differently, i.e., without an input depth image.
+The image transformer $E^{RGB}$ has a Vision Transformer (ViT) architecture [dosovitskiy2020image] with $16\times 16$ patch embeddings.
 The point transformer $E^{XYZ}$ uses a self-attention-based patch embedding design that differentiates between seen and unseen pixel points.
@@
 The decoder takes as input the encoder output, $R$, as well as a set of query points $Q\in\mathbb{R}^{N_{q}\times 3}$ (we choose $N_{q}=1024$).
-These inputs are passed to a transformer with architecture adopted from MCC and inspired by Masked Autoencoders (MAE) [26].
+These inputs are passed to a transformer with architecture adopted from MCC and inspired by Masked Autoencoders (MAE) [he2022masked].
 Each query token is passed to three output heads that infer binary occupancy $\sigma$, RGB color $c$, and hand-object segmentation $m$, such that:
@@
 How might we find a way to exploit these other models to our end?
-Specifically, recognition of objects is very well developed [51], and large collections of 3D object models [10, 11] exist and have facilitated advancements in 3D generative modeling [45, 57, 7, 43].
-Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [51] and a text-to-3D generative model [45] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image.
-We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [35].
+Specifically, recognition of objects is very well developed [gpt4v], and large collections of 3D object models [deitke2023objaverse, deitke2024objaverse] exist and have facilitated advancements in 3D generative modeling [luma, poole2022dreamfusion, chen2024text, long2024wonder3d].
+Thus, in order to further improve upon network-inferred object geometry, we exploit these “big data” models by using GPT-4(V) [gpt4v] and a text-to-3D generative model [luma] to automatically detect and “retrieve” a 3D object corresponding to the hand-held object in an image.
+We call this technique Retrieval-Augmented Reconstruction (RAR), somewhat in analogy to Retrieval-Augmented Generation in natural language processing [lewis2020retrieval].
 Given a representative frame of the video, we prompt GPT-4(V) to provide a detailed description of the object the hand is holding.
-This text caption is passed to Genie, a text-to-3D model from Luma [45], to retrieve realistic 3D object geometry and appearance (Genie is also used in [36]).
+This text caption is passed to Genie, a text-to-3D model from Luma [luma], to retrieve realistic 3D object geometry and appearance (Genie is also used in [li2023object]).
 Since Genie outputs geometry normalized to fit inside a unit cube, the scale of the 3D model must be adjusted with respect to the input 3D hand.
@@
 In order to (1) guarantee a globally optimal solution while (2) maintaining fixed compute time, we discretize the space of all possible parameter values to construct finite state spaces $\{\mathcal{R},\mathcal{T}\}$.
-The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [33].
+The 3D rotation group, SO(3), is discretized using the unit cube subdivision strategy proposed in [kurz2017discretization].
 Translation is defined as an offset from the mean location of the MCC-HO point cloud $\mu(X)$, and the space is discretized by constructing a small voxel grid centered at $\mu(X)$.
@@
 MCC-HO Chamfer Distance.
-The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric as defined in Equation 8.
+The Chamfer Distance (CD) between the MCC-HO point cloud and a transformed object model is the usual 3D evaluation metric.
 Note that when optimizing for rotation, the object model (centered at the origin) is first rotated and then translated to the mean location of the MCC-HO point cloud.
@@
 The MCC-HO point cloud gives us a reasonable estimate for object pose, but using this signal alone may not result in poses that align well with the input image (particularly for geometry with symmetries and in-the-wild objects).
-Thus, we additionally use DINOv2 [52] features to determine visual similarities between a transformed object model and the input image.
-For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [64].
+Thus, we additionally use DINOv2 [oquab2023dinov2] features to determine visual similarities between a transformed object model and the input image.
+For each rotation/translation state, the transformed object mesh is rendered using Pytorch3D [ravi2020accelerating].
 A PCA basis is constructed using the DINOv2 features of the input image, and the first three PCA components are computed for both the rendered features $\mathcal{F}_{j}$ and input image features $\mathcal{F}_{0}$.
 We compute the cosine similarity between the two as
-${\textstyle E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)}$ (6)
+$E_{DINO}=1-\frac{1}{2}\left(\frac{\text{PCA}(\mathcal{F}_{j})\cdot\text{PCA}(\mathcal{F}_{0})}{\max(||\text{PCA}(\mathcal{F}_{j})||_{2}\cdot||\text{PCA}(\mathcal{F}_{0})||_{2}),\epsilon)}+1\right)$} (6)
 where $\mathcal{F}_{j}$ and $\mathcal{F}_{0}$ are both masked by the input image object silhouette.
-When ground truth object masks are not available, we use SAM 2 [65].
-See Figure 3 for rotation results using example images from 100DOH [71].
+When ground truth object masks are not available, we use SAM 2 [ravi2024sam].
+See Figure 3 for rotation results using example images from 100DOH [shan2020understanding].
 Figure 3: We compute a PCA basis of DINOv2 features using all frames masked by the object silhouettes (one frame and its first three PCA components are shown on the left side of each example).
@@
 Given a sequence of video frames with timesteps $t$, we jointly optimize for per-frame rotations $R(t)$ and subsequently per-frame translations $T(t)$.
-The typical approach for temporal alignment is to use gradient-based optimization  [55, 95], which is prone to issues with local minima and has unbounded computation time.
-Aiming for robustness and efficiency, we propose using the Viterbi algorithm [80, 15, 29] to solve for a global solution to rigid object alignment across all frames of a video sequence.
+The typical approach for temporal alignment is to use gradient-based optimization  [patel2022learning, ye2023diffusion], which is prone to issues with local minima and has unbounded computation time.
+Aiming for robustness and efficiency, we propose using the Viterbi algorithm [viterbi1967error, forney1973viterbi, jurafsky2000speech] to solve for a global solution to rigid object alignment across all frames of a video sequence.
 The Viterbi algorithm is a dynamic programming approach to obtaining the maximum a posteriori probability of the most likely sequence of hidden states $Q=q_{1},q_{2},\ldots,q_{N}$ that results in a sequence of observed events $O=o_{1},o_{2},\ldots,o_{N}$.
@@
 Each cell of the Viterbi trellis, $v_{t}(j)$, is the probability of ending up in state $j$ after seeing the first $t$ observations and passing through the most probable state sequence $q_{1},\ldots,q_{t-1}$.
-Given the probability of being in every state at time $t-1$, the Viterbi probabilities are computed via:
+Given the probability of being in every state at time t1t-1, the Viterbi probabilities are computed via:
 $v_{t}(j)=\max_{i=1}^{|\mathcal{Q}|}\hskip 2.84526ptv_{t-1}(i)a_{ij}b_{j}(o_{t}),$ (7)
@@
 Section 7.2 defines the evaluation metrics used for comparisons.
-We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [4, 2, 41].
-Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [71] to demonstrate the feasibility and scalability of our automated model retrieval and fitting.
-Section 7.5 includes ablation studies for MCC-HO, RAR, and each component of our Viterbi temporal alignment algorithm.
+We also quantitatively (Section 7.3) and qualitatively (Section 7.4) evaluate against existing work and find that our approach achieves state-of-the-art performance on all three datasets used [chao2021dexycb, cao2021reconstructing, liu2022hoi4d].
+Also in Section 7.4.1, we provide qualitative results for our Retrieval-Augmented Reconstruction (RAR) approach applied to in-the-wild Internet videos from 100DOH [shan2020understanding] to demonstrate the feasibility and scalability of our automated model retrieval and fitting.
+Section 7.5 includes ablation studies for MCC-HO, improvements attributed to RAR, and the effect of conditioning hand-held object reconstruction on 3D hands.
 DexYCB MOW HOI4D
-F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
-HO [23] 0.24 0.48 4.76 0.03 0.06 49.8 0.28 0.51 3.86
-IHOI [94] - - - 0.13 0.24 23.1 0.42 0.70 2.7
-HORSE [58] 0.20 0.41 7.16 0.08 0.20 27.2 0.21 0.41 7.68
-MCC-HO (Ours) 0.36 0.60 3.74 0.15 0.31 15.2 0.52 0.78 1.36
-Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D. Chamfer Distance (cm2) and F-score (5mm, 10mm) are reported.
+F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) Vol ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) Vol ($\downarrow$) F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$) Vol ($\downarrow$)
+HO [hasson2019learning] 0.24 0.48 4.76 11.8 0.03 0.06 49.8 78.2 0.28 0.51 3.86 12.7
+IHOI [ye2022s] - - - - 0.13 0.24 23.1 20.4 0.42 0.70 2.7 0.91
+HORSE [Prakash2024HOI] 0.23 0.42 6.97 3.63 0.11 0.23 24.5 17.0 0.26 0.45 6.69 1.39
+MCC-HO (Ours) 0.36 0.60 3.74 2.42 0.15 0.31 15.2 13.8 0.52 0.78 1.36 1.94
+Table 1: We compare our method, MCC-HO, to prior works on held-out test images from DexYCB, MOW, and HOI4D.
+Chamfer Distance (cm2), F-score (5mm, 10mm), and Intersection Volume using 0.5cm voxels (as in HO [hasson2019learning], IHOI [ye2022s]) are reported.
 Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
-DiffHOI [89] 0.62 0.91 0.8
-G-HOP [96] 0.76 0.97 0.4
-MCC-HO + RAR (Ours) 0.91 0.98 0.28
-Table 2: Quantitative evaluation of video-based hand-held object reconstruction methods on the HOI4D test dataset. Mean metrics are reported for consistency with prior work.
+G-HOP [ye2024g] 0.61 0.89 0.8
+MCC-HO + RAR 0.74 0.91 0.64
+Table 2: Quantitative comparison to a state-of-the-art video-based hand-held object reconstruction method using the HOI4D test dataset. Mean metrics are reported for consistency with G-HOP.
 7.1 Datasets
@@
 HOI4D.
-The HOI4D dataset [41] consists of 4,000 lab videos of hands interacting with specified categories of common objects.
+The HOI4D dataset [liu2022hoi4d] consists of 4,000 lab videos of hands interacting with specified categories of common objects.
 We selected all sequences containing one of 6 hand-held object categories, and divide the sequences into training, validation and test sets.
-In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [95].
+In total, we used 1142 training sequences, 6 validation sequences, and 12 test sequences consistent with DiffHOI [ye2023diffusion].
 For the validation set, we selected one video per category and favored object instances that appear the least frequently in the dataset.
-DexYCB. The DexYCB dataset [4] contains 1,000 lab videos of hands interacting with 21 YCB objects. We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
+DexYCB.
+The DexYCB dataset [chao2021dexycb] contains 1,000 lab videos of hands interacting with 21 YCB objects.
+We used the “s0” dataset split, which contains 800 training, 40 validation, and 160 test sequences.
 MOW/RHOV.
-MOW [2] is a dataset of 512 images from 100 Days of Hands [71] labeled with 3D hands and objects.
-We use the same splits as [94] with 350 training images and 92 test images.
-Additionally, we use 110 videos from RHOV [55], each of which is a 2 second clip centered at the corresponding MOW frame.
+MOW [cao2021reconstructing] is a dataset of 512 images from 100 Days of Hands [shan2020understanding] labeled with 3D hands and objects.
+We use the same splits as [ye2022s] with 350 training images and 92 test images.
+Additionally, we use 110 videos from RHOV [patel2022learning], each of which is a 2 second clip centered at the corresponding MOW frame.
 In the RHOV dataset, 19 videos overlap with the MOW test set and thus we do not use them during training.
 100DOH.
-The 100 Days of Hands dataset [71] contains 131 days of footage collected from Internet videos showing hands interacting with objects.
+The 100 Days of Hands dataset [shan2020understanding] contains 131 days of footage collected from Internet videos showing hands interacting with objects.
 The vast majority of videos do not have associated 3D hand or object labels, so we evaluate how well our approach generalizes using examples from this unlabeled subset.
-For DexYCB [4], MOW [2], and HOI4D [41], all results generated using our approach use the ground truth 3D hand provided by each dataset.
-For 100DOH [71] results (e.g., examples not contained in MOW), we use HaMeR [56] to estimate the 3D hand.
+For DexYCB [chao2021dexycb], MOW [cao2021reconstructing], and HOI4D [liu2022hoi4d], all results generated using our approach use the ground truth 3D hand provided by each dataset.
+For 100DOH [shan2020understanding] results (e.g., examples not contained in MOW), we use HaMeR [pavlakos2023reconstructing] to estimate the 3D hand.
 7.2 Evaluation Metrics
@@
 Both metrics are computed on point sets.
-The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [69] and HOI4D optimizes hand pose using 2D keypoint annotations.
-For each example, we sample 10k points on the surface of the ground truth mesh consistent with [94, 95] (no transformations are applied to the ground truth geometry).
+The 3D hand provided by each dataset is used as input; note that MOW uses FrankMocap [rong2020frankmocap] and HOI4D optimizes hand pose using 2D keypoint annotations.
+For each example, we sample 10k points on the surface of the ground truth mesh consistent with [ye2022s, ye2023diffusion] (no transformations are applied to the ground truth geometry).
 The MCC-HO predicted point clouds are supersampled/subsampled to obtain 10k points as well.
-In order to account for unknown ground truth camera parameters, ICP with scaling [64] is used to align all predicted geometry with the ground truth mesh.
+In order to account for unknown ground truth camera parameters, ICP with scaling [ravi2020accelerating] is used to align all predicted geometry with the ground truth mesh.
 Median metrics are reported unless otherwise noted.
-Chamfer Distance. The CD metric we use between two point sets $A,B\in\mathbb{R}^{N\times 3}$ is:
-$CD(A,B)=\frac{1}{N}\sum_{a\in A}\min_{b\in B}||a-b||^{2}_{2}+\frac{1}{N}\sum_{b\in B}\min_{a\in A}||a-b||^{2}_{2},$ (8)
-where the point sets should have the same size. This equation is consistent with prior work [5, 6, 94, 95]. We report CD in cm2 as in[5, 6].
-F-score. The F-score at a given threshold $t$ is defined as a function of precision $P$ and recall $R$:
-$F_{1}(X_{pd},X_{gt})=2\vspace{1mm}\frac{P(X_{pd},X_{gt},t)\cdot R(X_{pd},X_{gt},t)}{P(X_{pd},X_{gt},t)+R(X_{pd},X_{gt},t)}.$ (9)
 7.3 Quantitative Evaluation
-We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [23, 94, 58].
+We quantitatively evaluate MCC-HO on a variety of labeled hand-object datasets and in comparison to existing model-free reconstruction approaches [hasson2019learning, ye2022s, Prakash2024HOI].
 See Table 1.
@@
 For DexYCB and HOI4D, we use every 5th frame of each test video sequence.
-The results for IHOI [94] and HO [23] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [95] papers; otherwise, we used available public repositories of prior work.
-Because AlignSDF [5] and gSDF [6] trained DexYCB-specific models, whereas we train a single model on multiple datasets, we do not compare to their published metrics.
+The results for IHOI [ye2022s] and HO [hasson2019learning] evaluated on the MOW dataset are obtained from the IHOI and DiffHOI [ye2023diffusion] papers; otherwise, we used available public repositories of prior work.
 Our method achieves state-of-the-art performance for the Chamfer Distance and F-score metrics on all three datasets, implying that our technique is not overfitted to any particular dataset.
 We conjecture that increasing the number of unique object instances seen during training is important to model generalization.
-In addition, the pretrained MCC model provides a strong object prior for initialization.
-We also compare to existing video-based approaches for RGB hand-held object reconstruction in Table 2, noting that the ground truth object is used for our RAR rigid alignment (a reasonable comparison as DiffHOI [95] and GHOP [96] train category-specific diffusion models using the ground truth HOI4D 3D objects).
+Thus, the pretrained MCC model provides a strong object prior for initialization.
+We also compare to G-HOP [ye2024g], a state-of-the-art video-based approach for RGB hand-held object reconstruction in Table 2.
+Our method (using generated Genie objects for RAR) outperforms the generalist G-HOP model (that was trained on seven hand-object datasets [brahmbhatt2020contactpose, chao2021dexycb, corona2020ganhand, liu2022hoi4d, taheri2020grab, yang2022oakink]) in all three metrics.
 7.4 Qualitative Evaluation
 We qualitatively evaluate (1) the reconstruction quality of MCC-HO and (2) how well retrieved object models using RAR can be rigidly aligned to the predicted point clouds in 3D (a necessary and sufficient condition for downstream tasks).
-Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of rigidly aligning the ground truth CAD model with each point cloud.
+Figure 4 shows predicted point clouds for a number of MOW test examples, as well as the result of RAR.
 Since MCC-HO is trained with monocular images, the predicted RGB colors are only valid when viewed from the given camera.
@@
 Additionally, the object in contact with the hand is correctly identified by MCC-HO without input segmentation maps.
+Figure 5 shows qualitative comparisons to existing work, where the combination of MCC-HO + RAR allows us to obtain more realistic object geometry.
 Figure 4: MCC-HO results for MOW test examples.
-The input image (top, left), network-inferred hand-object point cloud (top, right), rigid alignment of the ground truth mesh with the point cloud (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
-Figure 5: Our approach applied to images in the 100 Days of Hands dataset [71].
-For each image, we estimate 3D hands using HaMeR [56], inference object geometry using MCC-HO, retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie), and rigidly align the 3D model with the inferred point cloud from MCC-HO.
+The input image (top, left), network-inferred hand-object point cloud (top, right), RAR (bottom, left), and an alternative view of the point cloud (bottom, right) are shown.
+Figure 5: Qualitative comparisons using MOW test examples.
+*Note that the MCC-HO point clouds are rendered as a mesh via Poisson surface reconstruction, which can introduce artifacts not attributed to our method.
+The last column is MCC-HO + RAR.
+Figure 6: Our approach applied to in-the-wild images from the 100 Days of Hands dataset [shan2020understanding].
+For each image, we estimate 3D hands using HaMeR [pavlakos2023reconstructing], inference object geometry using MCC-HO and retrieve a 3D model using RAR (e.g., GPT-4(V) and Genie).
 7.4.1 Retrieval-Augmented Reconstruction
 The ultimate goal of RAR is to provide a scalable paradigm for creating large hand-object 3D datasets; to this end, we explore the potential to scale up our approach on unlabeled images and videos.
-Figure 5 shows our results for 9 images from 100DOH [71] that do not have 3D ground truth (e.g., not in MOW).
-Estimated 3D hands for each image are obtained using HaMeR [56].
-At test time, our method does not require an input hand-object segmentation mask [2, 55], depth image [89], or hand joint locations [94, 8].
+Figure 6 shows our results for 9 images from 100DOH [shan2020understanding] that do not have 3D ground truth (e.g., not in MOW).
+Estimated 3D hands for each image are obtained using HaMeR [pavlakos2023reconstructing].
+At test time, our method does not require an input hand-object segmentation mask [cao2021reconstructing, patel2022learning], depth image [wu2023multiview], or hand joint locations [ye2022s, choi2023handnerf].
 Please see the supplementary materials for additional results.
@@
 MCC Baseline.
-Because our model is built on top of MCC [89], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction.
-For MCC, we use the released model that was trained on the CO3Dv2 dataset [67] and apply an existing monocular depth estimator [63] to obtain depth maps (as MCC takes an RGB-D image as input).
+Because our model is built on top of MCC [wu2023multiview], we begin by evaluating whether MCC-HO leads to performance gains for the task of hand-held object reconstruction.
+For MCC, we use the released model that was trained on the CO3Dv2 dataset [reizenstein21co3d] and apply an existing monocular depth estimator [Ranftl2021] to obtain depth maps (as MCC takes an RGB-D image as input).
 Table 3 compares 3D metrics between MCC and MCC-HO computed on the MOW test set.
@@
 Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
-MCC [89] 0.10 0.21 31.6
+MCC [wu2023multiview] 0.10 0.21 31.6
+MCC-HO 0.15 0.31 15.2
+Table 3: Analysis of MCC-HO performance gains over our object reconstruction baseline, MCC, using the MOW test dataset.
+Effectiveness of Input Hand.
+Table 4 analyzes the effectiveness of hand pose prediction input by evaluating MCC-HO with and without the input hand mesh.
+As we expect, the model performs significantly worse without the input hand geometry.
+The experimental results further corroborate the hypothesis that grounding the 3D object prediction with respect to the hand coordinate system significantly constrains the reconstruction problem.
+Model Parameters F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
+MCC [wu2023multiview] 0.10 0.21 31.6
+MCC-HO (w/o hand) 0.03 0.12 71.6
 MCC-HO 0.15 0.31 15.2
 Table 4: Analysis of how hand-held object reconstruction benefits from conditioning on the hand using the MOW test dataset.
-RAR Success Rate.
-In order to quantify the success rate of RAR (e.g., whether a generated 3D object matches the input image), we use GPT-4(V) and Genie to automatically generate objects for all 512 images in the MOW dataset.
-Genie is able to generate an object 82% of the time, and otherwise returned an error.
-The results are included in the supplementary materials.
-Video Alignment Components.
-We aim to understand the contributions of each cost term in our Viterbi algorithm, e.g., MCC-HO and DINOv2.
-See Table 4.
-In order to best assess the quality of rigid alignment, the ICP postprocess described in Section 7.2 is not applied.
-The results demonstrate that both terms contribute to the efficacy of our Viterbi algorithm; in practice, we observe that using both 2D and 3D cues is particularly important for objects that lie outside the training distribution of MCC-HO (DexYCB consists of commonplace YCB objects).
-Viterbi Cost Terms F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
-MCC-HO 0.26 0.52 4.18
-DINOv2 0.23 0.45 6.38
-MCC-HO + DINOv2 0.27 0.52 4.10
-Table 4: Analysis of how each of our Viterbi cost terms contributes to the final temporal alignment on the DexYCB test dataset.
+RAR Reconstruction Improvements.
+RAR leverages text-to-3D generative models to synthesize realistic geometry, and we quantitatively evaluate whether using RAR in conjunction with MCC-HO leads to improved results.
+Table 5 demonstrates that RAR (with Genie objects) can indeed help improve the object reconstruction quality of monocular reconstruction methods such as MCC-HO.
+Method F-5 ($\uparrow$) F-10 ($\uparrow$) CD ($\downarrow$)
+MCC-HO 0.15 0.31 15.2
+MCC-HO + RAR 0.16 0.44 9.41
+Table 5: Analysis of how RAR improves object reconstruction. Median Chamfer Distance (cm2) and F-score (5mm, 10mm) metrics are computed on the MOW test dataset.
 8 Conclusion
@@
 We additionally present Retrieval-Augmented Reconstruction (RAR), an automatic method for object model retrieval that leverages recent advances in large language/vision models and 3D object datasets.
-These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [71].
+These two approaches can be combined to scale the amount of labeled hand-object interaction data, as suggested by our results using unlabeled 100DOH videos [shan2020understanding].
 Limitations.